### v2

# Stable cohomology of universal character varieties

Abstract.
We study the universal $\operatorname{PGL}_{n}$-character variety over $\mathcal{M}_{g}$ whose fiber over a point $[C]$ is the space of $\operatorname{PGL}_{n}$-local systems on the curve $C$. We use nonabelian Hodge theory and properties of Saito’s mixed Hodge modules to show that the Leray-Serre spectral sequence for the projection to $\mathcal{M}_{g}$ degenerates at $E_{2}$. As an application, we prove that the rational cohomology of these varieties stabilizes in the limit as $g\to\infty$ and compute the stable limit. We also deduce similar results for the universal $G$-character variety over $\mathcal{M}_{g,1}$ whose fiber over a punctured curve is the variety of $G$-local systems with fixed central monodromy around the puncture, for $G=\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$.
1. Introduction
The degree $k$ cohomology group of the moduli space of curves of genus $g$ is independent of $g$ for $k\leq\frac{2g-2}{3}$. These stable cohomology groups were computed by Madsen and Weiss [31], confirming a conjecture of Mumford [39]. Subsequently, Ellenberg, Venkatesh, and Westerland initiated an analogous study of cohomological stability for Hurwitz spaces, which parametrize branched $G$-covers of $\mathbb{A}^{1}$ for a finite group $G$ or, equivalently, $G$-local systems on $\mathbb{A}^{1}$ minus finitely many branch points [14]. More recently, Landesman and Levy computed the stable cohomology groups of these $G$-Hurwitz spaces [29] as the number of branch points goes to $\infty$, and Putman [43, Theorem C] proved a homological stability result for $G$-Hurwitz spaces of a genus $g$ curve with one boundary component in the limit as $g\to\infty$.
Inspired by this recent progress, we investigate analogous stability phenomena for the cohomology groups of universal character varieties that parametrize $G$-local systems on smooth curves of genus $g$, for certain algebraic groups $G$. We prove that these groups stabilize in the limit as $g\to\infty$ and compute the stable limits.
Our first such result is for $G=\operatorname{PGL}_{n}$. We consider the universal character variety $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n)\to\mathcal{M}_{g}$ whose fiber over $[C]$ is the character variety
$\mathrm{Hom}(\pi_{1}(C),\operatorname{PGL}_{n}(\mathbb{C}))/\operatorname{PGL}_{n}(\mathbb{C})],$
parametrizing $\operatorname{PGL}_{n}(\mathbb{C})$-local systems on $C$ via the monodromy representation. This character variety breaks into pieces indexed by $\mathbb{Z}/n\mathbb{Z}$, based on the obstruction to lifting a homomorphism to $\operatorname{PGL}_{n}(\mathbb{C})$ to $\operatorname{GL}_{n}(\mathbb{C})$, which is locally constant. We study the stable cohomology of the components $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ corresponding to $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, which are smooth; the appendix explains why this is the same as stable intersection cohomology for any $d\in(\mathbb{Z}/n\mathbb{Z})$. See §2 for precise definitions and further details.
To describe the stable cohomology groups we use the following notation. Let
$V_{g}:=H^{1}(S_{g};\mathbb{Q})$ (1)
be the first rational cohomology group of a closed orientable surface of genus $g$, viewed as a $\operatorname{Mod}_{g}$-module or, equivalently, a local system on $\mathcal{M}_{g}$. We consider $V_{g}$ as a graded object, supported in degree $1$, so $V_{g}[-i]$ is supported in degree $i+1$. By convention, the grading on the cohomology of a graded local system is given by the sum of the cohomological degree and the degree of the local system, e.g. $H^{k}(\mathcal{M}_{g};V_{g}[-2])$ is of degree $k+3$. We also define
$V_{\circ}:=\mathbb{Q}\oplus\mathbb{Q}[-2],\quad\quad W_{n}:=\mathbb{Q}\oplus\mathbb{Q}[-2]\oplus\cdots\oplus\mathbb{Q}[2-2n],\quad\quad\quad\widetilde{W}_{n}:=W_{n}/\mathbb{Q}.$ (2)
Here, the graded vector space $V_{\circ}$ may be thought of as the even cohomology of a closed orientable surface of arbitrary genus, supported in degrees 0 and 2. Likewise, $W_{n}$ and $\widetilde{W}_{n}$ may be viewed as the cohomology of $\mathbb{P}^{n-1}$ and the reduced cohomology of $\mathbb{P}^{n-1}$, respectively.
Theorem 1.1.
For $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d);\mathbb{Q})$ is isomorphic to the degree $k$ part of the graded vector space
$H^{*}\big(\mathcal{M}_{g};\textstyle\bigwedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n}).$ (3)
Corollary 1.2.
For $k\leq\frac{2g-2}{3}$, the group $H^{k}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d);\mathbb{Q})$ does not depend on $g$ and vanishes if $k$ is odd. If, in addition, $k\leq 2n-1$ then this group is independent of $n$.
Corollary 1.2 follows from Theorem 1.1 via Looijenga’s results on the cohomology of $\mathcal{M}_{g}$ with symplectic coefficients [30]. See Section 4.2 for further details on the stable cohomology groups. The dimensions of the stable cohomology groups of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,1)$ for $n\leq 7$ in even degrees up to 20 are displayed in Figure 1.
The following statement is a key technical step in the proof of Theorem 1.1, which is also of independent interest. Let $\pi\colon\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ be the forgetful map taking a $\operatorname{PGL}_{n}$-local system on a smooth curve to the underlying curve.
Theorem 1.3.
Let $d$ and $n$ be coprime positive integers, and let $g\geq 2$. The Leray–Serre spectral sequence for $\pi\colon\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ degenerates at $E_{2}$.
We prove this degeneration using the relative nonabelian Hodge correspondence [49], purity of the cohomology of the moduli space of Higgs bundles [18, Corollary 1.3.2], and a splitting result for pure Hodge modules [45, (4.5.4)]. We also present a second proof using only classical Hodge theory and manipulations in the derived category of constructible sheaves. Theorem 1.1 is deduced from the degeneration of the Leray-Serre spectral sequence for $\pi$ using previously known results on generators and relations of the cohomology of the fibers ([32, Theorem 7] and [19, Lemma 4.1.12]).
We also consider universal $G$-character varieties over $\mathcal{M}_{g,1}$, for $G$ = $\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$. The fiber of the universal $G$-character variety $M^{G}_{B}(\mathcal{C}_{g,1},n,d)$ over a punctured curve $C$ is given by
$\{\phi\in\mathrm{Hom}(\pi_{1}(C),G(\mathbb{C})):\phi(\gamma)=e^{2\pi id/n}\operatorname{Id}\}/\operatorname{PGL}(\mathbb{C}).$
Here, $\gamma$ is a chosen loop around the puncture. These universal $G$-character varieties admit natural quotient maps to the pullback of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$, and the cohomology groups of the character varieties for these three groups are closely related.
Theorem 1.4.
For $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d);\mathbb{Q})$ is isomorphic to the degree $k$ part of the graded vector space
$H^{*}(\mathcal{M}_{g,1},\textstyle{\bigwedge^{\bullet}}(V_{g}\otimes\widetilde{W}_{n}))\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n}).$
Similarly, for $k\leq 2(g-1)(n-1)+2$, the cohomology group $H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d);\mathbb{Q})$ is isomorphic to the degree $k$ part of the graded vector space
$H^{*}(\mathcal{M}_{g,1},\textstyle{\bigwedge^{\bullet}}(V_{g}\otimes W_{n}))\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n}).$
For $k\leq(2g-2)/3$, these groups are independent of $g$ and vanish when $k$ is odd.
The final statements about stable cohomology in low degrees follows, just as Corollary 1.2 follows from Theorem 3.1, using results of Kawazumi and Randal-Williams on the cohomology of $\mathcal{M}_{g,1}$ with twisted coefficients [26, 44]. See Section 4.3 for further discussion of the stable cohomology. The dimensions of the stable cohomology groups of $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},2,1)$ in even degrees up to 20 are displayed in Figure 1.
Remark 1.5.
Although we initially hoped that each universal degree $d$ coprime to $n$ character variety could naturally be defined over $\mathcal{M}_{g}$ instead of $\mathcal{M}_{g,1}$, we no longer believe that this is the case for $\operatorname{GL}$ and $\operatorname{SL}$; see Remark 2.4. This is also consistent with our belief that the most natural path to proving a relative nonabelian Hodge correspondence in coprime degree would be to reinterpret the question in terms of terms of degree 0 filtered objects on a punctured curve, extending Simpson’s proof of a fiberwise bijection [48]. However, to our knowledge such a result is not in the literature, and so we reduce to using Simpson’s degree 0 relative nonabelian Hodge result (for $\operatorname{PGL}$) [49].
Remark 1.6.
By Theorem 1.1, the rational cohomology of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. We do not know any way of deducing this from first principles. The fibers of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ are naturally viewed as algebraic varieties defined over $\mathbb{Q}[e^{2\pi i/n}]$. The fibers for varying $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ are Galois conjugate, giving an isomorphism of cohomology (with $\mathbb{Q}_{\ell}$-coefficients). Such arguments do not apply to the total space of the universal character varieties, which are constructed analytically (Definition 2.1). Our proof of Theorem 1.1 depends on the relative non-abelian Hodge correspondence, which produces a homeomorphism over $\mathcal{M}_{g}$ between the universal character variety $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Betti side and a universal moduli space of semistable Higgs bundles of degree $d$ $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ on the Dolbeault side. The fact that the cohomology is independent of the choice of $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ does not have an obvious geometric explanation in the Higgs bundle setting, though the result fits into a more general conjecture about the behavior of Gopakumar-Vafa invariants on Calabi-Yau 3-folds. See [34, §0.3].
$2k$ 0 2 4 6 8 10 12 14 16 18 20
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},2,1);\mathbb{Q}\big)$ 1 2 5 11 23 45 87 160 290 512 889
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},3,1);\mathbb{Q}\big)$ 1 2 6 14 33 71 152 307 612 1181 2243
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},4,1);\mathbb{Q}\big)$ 1 2 6 15 36 81 180 380 788 1588 3138
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},5,1);\mathbb{Q}\big)$ 1 2 6 15 37 84 190 408 863 1772 3574
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},6,1);\mathbb{Q}\big)$ 1 2 6 15 37 85 193 418 891 1847 3760
$\dim_{\mathbb{Q}}H^{2k}\big(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},7,1);\mathbb{Q}\big)$ 1 2 6 15 37 85 194 421 901 1875 3835
$\dim_{\mathbb{Q}}H^{2k}\big(M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},2,1);\mathbb{Q}\big)$ 1 3 9 22 51 109 225 443 849 1579 2874
Figure 1. Betti numbers of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,1)$ for $2\leq n\leq 7$ and $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},2,1)$, for even degrees $2k\leq 20$ in the stable range $2k\leq(2g-2)/3$.
Acknowledgments. This work grew out of discussions from the AIM workshop on “Motives and mapping class groups.” We thank AIM for the stimulating and supportive environment and our fellow workshop participants for many engaging conversations. We are especially grateful to Aaron Landesman for proposing this project and Eduard Looijenga for invaluable input in its early stages. We also thank Bhargav Bhatt and Davesh Maulik for helpful conversations related to this work, and Daniel Litt for comments on an earlier version of this paper. A.L. and F.J. were partially supported by an NSF Graduate Research Fellowship under grants no. 2141064 and no. 2140001 respectively. S.P. was partially supported by NSF grant DMS–2542134 and carried out portions of this work while visiting the Institute for Advanced Study, supported by the Charles Simonyi Endowment. X.Z. is grateful to Max Planck Institute for Mathematics in Bonn for its hospitality.
2. Universal character varieties
In this section, we start by defining the universal $\operatorname{PGL}_{n}$-character variety over $\mathcal{M}_{g}$. We then define the universal $G$-character varieties over $\mathcal{M}_{g,1}$, for $G$ equal to $\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$. We also discuss the natural relations between these spaces.
Let $S_{g}$ be a closed oriented surface of genus $g\geq 2$, and let $\operatorname{Mod}_{g}:=\pi_{0}(\operatorname{Diff}^{+}(S_{g}))$ be its mapping class group. The group $\operatorname{PGL}_{n}(\mathbb{C})$ acts by conjugation on the affine variety
$X_{g,n}:=\mathrm{Hom}(\pi_{1}(S_{g}),\operatorname{PGL}_{n}(\mathbb{C})).$
Let $\mu_{n}\subset\mathbb{C}$ be the group of $n$th roots of unity, which we identify with $\mathbb{Z}/n\mathbb{Z}$ via $d\mapsto e^{2\pi id/n}$. By the short exact sequence
$1\to\mu_{n}\to\operatorname{SL}_{n}(\mathbb{C})\to\operatorname{PGL}_{n}(\mathbb{C})\to 1,$
the obstruction to lifting a representation $\phi\colon\pi_{1}(S_{g})\to\operatorname{PGL}_{n}(\mathbb{C})$ to $\operatorname{SL}_{n}(\mathbb{C})$ is given by an element of $H^{2}(S_{g};\mu_{n})=\mathbb{Z}/n\mathbb{Z}$, and this obstruction is locally constant on $X_{g,n}$. We write $X_{g,n,d}$ for the subvariety parametrizing homomorphisms with obstruction class $d\in\mathbb{Z}/n\mathbb{Z}$. Via the Riemann-Hilbert correspondence, $X_{g,n,d}$ parametrizes framed $\operatorname{PGL}_{n}$-bundles with flat connection that lift to $C^{\infty}$-vector bundles of degree $d$ mod $n$.
We now define the degree $d$ $\operatorname{PGL}_{n}$-character variety to be the quotient stack
$\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d):=[X_{g,n,d}/\operatorname{PGL}_{n}(\mathbb{C})],$
where $\operatorname{PGL}_{n}(\mathbb{C})$ acts by conjugation on the target. Likewise, we define the degree $d$ coarse $\operatorname{PGL}_{n}$-character variety to be the affine GIT quotient
$M_{B}^{\operatorname{PGL}}(S_{g},n,d)=X_{g,n,d}/\!\!/\operatorname{PGL}_{n}(\mathbb{C}).$ (4)
In other words, $M_{B}^{\operatorname{PGL}}(S_{g},n,d)$ is the affine variety whose coordinate ring is the ring of $\operatorname{PGL}_{n}(\mathbb{C})$-invariant regular functions on $X_{g,n,d}$. The subscript $B$ indicates that these character varieties appear on the Betti side of the nonabelian Hodge correspondence.
Assume $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$. Then $X_{g,n,d}$ is smooth and irreducible, and $\operatorname{PGL}_{n}(\mathbb{C})$ acts with finite stabilizers [19, Theorem 2.2.12 and Corollary 3.5.5]. In particular, $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ is a smooth orbifold, and projection to the coarse space $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)\to M_{B}^{\operatorname{PGL}}(S_{g},n,d)$ induces an isomorphism
$H^{*}(M_{B}^{\operatorname{PGL}}(S_{g},n,d),\mathbb{Q})\cong H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d),\mathbb{Q}).$ (5)
Let $\mathcal{T}_{g}$ denote the Teichmüller space of $S_{g}$, which parametrizes complex structures on $S_{g}$ up to the action of the identity component of $\operatorname{Diff}^{+}(S_{g})$. It carries a natural action of $\operatorname{Mod}_{g}$, and the quotient is the moduli space of curves $\mathcal{T}_{g}/\operatorname{Mod}_{g}\cong\mathcal{M}_{g}$. The outer action of $\operatorname{Mod}_{g}$ on $\pi_{1}(S_{g})$ also induces a $\operatorname{Mod}_{g}$-action on the character variety $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$.
Definition 2.1.
For $d$ coprime to $n$, the universal degree $d$ $\operatorname{PGL}$-character variety is
$\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\coloneqq\mathcal{T}_{g}\times_{\operatorname{Mod}_{g}}\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d).$
By construction, this universal character variety is a smooth and irreducible orbifold. It is a smooth fiber bundle over the moduli stack $\mathcal{M}_{g}$ with fiber $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$.
Remark 2.2.
Although the base and fiber of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ each carry a natural algebraic structure, our construction does not immediately endow the total space with such a structure. In order to apply techniques from algebraic geometry to study the rational cohomology of $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$, we pass to a homeomorphic algebraic moduli space of semistable Higgs bundles on the Dolbeault side of the nonabelian Hodge correspondence.
We now consider degree $d$ $G$-character varieties for a surface with a puncture, for $G=\operatorname{SL}_{n}$ or $\operatorname{GL}_{n}$. Let $S_{g,1}$ be an orientable surface of genus $g$ with one puncture, and let $\gamma$ denote a loop around the puncture. For $d\in\mathbb{Z}/n\mathbb{Z}$, we consider the affine variety
$Y^{G}_{g,n,d}:=\{\phi\in\mathrm{Hom}(\pi_{1}(S_{g,1}),G(\mathbb{C})):\phi(\gamma)=e^{2\pi id/n}\operatorname{Id}_{n}\},$ (6)
on which $G(\mathbb{C})$ acts by conjugation. If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $Y^{G}_{g,n,d}$ is smooth and irreducible, and the $G(\mathbb{C})$-action factors through a free action of $\operatorname{PGL}_{n}(\mathbb{C})$ [19, Corollary 2.2.7]. In these cases, the degree $d$ $G$-character variety is the quotient
$M_{B}^{G}(S_{g,1},n,d)\coloneqq Y^{G}_{g,n,d}/\operatorname{PGL}_{n}(\mathbb{C}).$ (7)
Unlike $\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$, which is an orbifold (or smooth Deligne–Mumford stack with nontrivial stabilizers), the space $M_{B}^{G}(S_{g,1},n,d)$ is a manifold (or smooth algebraic variety).
Because the outer action of the mapping class group $\operatorname{Mod}_{g,1}\coloneq\pi_{0}(\operatorname{Diff}^{+}(S_{g,1}))$ on $\pi_{1}(S_{g,1})$ preserves the conjugacy class of $\gamma$ [15, Theorem 8.8], $\mathcal{M}_{B}^{G}(S_{g,1},n,d)$ inherits an action of the mapping class group $\operatorname{Mod}_{g,1}$ of $S_{g,1}$, by precomposition. Let $\mathcal{T}_{g,1}$ denote the Teichmüller space of $S_{g,1}$, which likewise carries a natural action of $\operatorname{Mod}_{g,1}$.
Definition 2.3.
The universal degree $d$ $G$-character variety for $G=\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$ is
$M_{B}^{G}(\mathcal{C}_{g,1},n,d)\coloneqq\mathcal{T}_{g,1}\times_{\operatorname{Mod}(S_{g,1})}M_{B}^{G}(S_{g,1},n,d).$ (8)
Remark 2.4.
When $d=0$, the condition $\phi(\gamma)=\operatorname{Id}_{n}$ from (6) is equivalent to the condition that $\phi$ factors through the fundamental group of the unpunctured surface $\pi_{1}(S_{g,1})\to\pi_{1}(S_{g}).$ The outer action of $\operatorname{Mod}_{g,1}$ on $\pi_{1}(S_{g})$ then factors through the outer action of $\operatorname{Mod}_{g}$, and $M_{B}^{G}(\mathcal{C}_{g,1},n,0)$ is the pullback of an analogously defined universal degree 0 $G$-character variety over $\mathcal{M}_{g}$. Hence, one can study degree 0 relative $G$-character varieties for families of unpunctured curves, as in [49, §6 and §9].
However, if $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$ is not the pullback of a space over $\mathcal{M}_{g}$, because the kernel of the map $\pi_{1}(\mathcal{M}_{g,1})\to\pi_{1}(\mathcal{M}_{g})$, acts nontrivially on $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$. This can be seen from the description of its outer action on $\pi_{1}(S_{g})$ in terms of Dehn twists [15, Fact 4.7].
We define a relative degree $d$ $\operatorname{PGL}$-character variety over $\mathcal{M}_{g,1}$ (or indeed any family of curves) by pullback:
$M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g,1},n,d):=\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\times_{\mathcal{M}_{g}}\mathcal{M}_{g,1}.$
This is analogous to the definition for $G=\operatorname{GL}_{n}$ or $\operatorname{SL}_{n}$, since the $\operatorname{PGL}$-character variety parametrizes homomorphisms that factor through $\pi_{1}(S_{g})$.
We now discuss relations among the universal $G$-character varieties for varying groups $G$. Note that $\Gamma:=\mathrm{Hom}(\pi_{1}(S_{g,1}),\mu_{n})$ acts on $Y_{g,n,d}^{G}$ by multiplication, via the inclusion of $\mu_{n}$ in the center of $G$, and this action descends to an action on the character variety $M_{B}^{G}(S_{g,1},n,d)$. Passing to the universal setting, the relative group object
$\widetilde{\Gamma}:=\mathcal{T}_{g,1}\times_{\operatorname{Mod}_{g,1}}\Gamma$
acts on $M_{B}^{G}(\mathcal{C}_{g,1},n,d)$ over $\mathcal{M}_{g,1}$.
Proposition 2.5.
The quotient of $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)$ by $\widetilde{\Gamma}$ over $\mathcal{M}_{g,1}$ is the morphism to $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g,1},n,d)$ induced by the projection $\operatorname{SL}_{n}(\mathbb{C})\to\operatorname{PGL}_{n}(\mathbb{C})$.
Similarly, the quotient of $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\times_{\mathcal{M}_{g,1}}M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)$ by $\widetilde{\Gamma}$ acting diagonally is the morphism to $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d)$ induced by the product $\mathbb{C}^{*}\times\operatorname{SL}_{n}(\mathbb{C})\to\operatorname{GL}_{n}(\mathbb{C})$.
Proof.
First, note that $\operatorname{PGL}_{n}=\operatorname{SL}_{n}/\mu_{n}$, where $\mu_{n}$ acts by its central embedding. Similarly, $\operatorname{GL}_{n}=(\operatorname{GL}_{1}\times\operatorname{SL}_{n})/\mu_{n}$, where $\mu_{n}$ acts diagonally. Taking $\mathrm{Hom}(\pi_{1}(S_{g,1}),-)$ gives
$X_{g,n,d}=Y_{g,n,d}^{\operatorname{SL}}/\Gamma\quad\mbox{ and }\quad Y^{\operatorname{GL}}_{g,n,d}=(Y_{g,1,0}^{\operatorname{GL}}\times Y_{g,n,d}^{\operatorname{SL}})/\Gamma.$
In each case, the action of $\mu_{n}$ commutes with both conjugation by $\operatorname{PGL}_{n}$ and also the action of $\operatorname{Mod}_{g,1}$ on $\pi_{1}(S_{g,1})$, inducing the required morphism of universal character varieties. ∎
3. Degeneration of the Leray spectral sequence
In this section, we prove the degeneration of the Leray spectral sequence for the universal degree $d$ $\operatorname{PGL}$-character variety $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$. Similar results for universal degree $d$ $\operatorname{GL}$- and $\operatorname{SL}$-character varieties follow as corollaries. Our proof uses the relative nonabelian Hodge correspondence for $\operatorname{PGL}$. This relates the universal degree $d$ $\operatorname{PGL}_{n}$-character variety on the Betti side of the correspondence to the universal moduli space of semistable degree $d$ $\operatorname{PGL}_{n}$-Higgs bundles on the Dolbeault side. On the Dolbeault side, the projection to $\mathcal{M}_{g}$ shares many cohomological properties of a smooth and proper morphism of algebraic varieties, despite not being proper. Using these properties, we prove the following strong form of Theorem 1.3.
Theorem 3.1.
Let $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and let $\pi\colon\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ be the projection from the universal degree $d$ $\operatorname{PGL}$-character variety. Then $R\pi_{*}\mathbb{Q}\cong\oplus_{i}R^{i}\pi_{*}\mathbb{Q}[-i]$ in the bounded derived category of constructible $\mathbb{Q}$-sheaves on $\mathcal{M}_{g}$. In particular, the Leray spectral sequence
$E_{2}^{pq}=H^{p}(\mathcal{M}_{g},R^{q}\pi_{*}\mathbb{Q})\implies H^{p+q}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{M}_{g},n,d);\mathbb{Q})$
degenerates at $E_{2}$.
Remark 3.2.
The first statement in Theorem 3.1 is a universal form of the degeneration of the Leray spectral sequence in the following sense: an object $A$ in a bounded derived category $D^{b}(\mathcal{A})$ splits as the direct sum of its cohomology sheaves if and only if for any cohomological functor $F$ from $D^{b}(\mathcal{A})$ to an abelian category, the corresponding spectral sequence $E_{2}^{pq}=F(\mathcal{H}^{q}(A)[p])\implies F(A[p+q])$ degenerates at $E_{2}$ [9, Proposition 1.2].
We present two proofs of Theorem 3.1. The first (§3.4) uses Saito’s theory of mixed Hodge modules. This is efficient, but relies on more substantial machinery than the second proof (§3.5), which involves only classical Hodge theory and some manipulations in the derived category of constructible sheaves.
Remark 3.3.
One may also give a third proof of Theorem 3.1, similar in spirit to the two presented here, but different in technical detail, via spreading out over a finitely generated $\mathbb{Z}$-algebra, reducing modulo large finite primes, and applying the theory of weights in the $\ell$-adic cohomology of local systems on varieties over finite fields, along the lines of [2, Theorem 5.4.5]. Like the two proofs presented here, this argument relies on the theory of weights in the cohomology of local systems on algebraic varieties [12]. However, to present this third proof in detail would take us too far from the setting of this paper, which is rooted in geometry over the complex numbers.
In §3.6, we deduce the following corollaries:
Corollary 3.4.
Let $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, and let $\pi\colon M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}$ be the projection map from the universal degree $d$ $\operatorname{GL}_{n}$-character variety. Then $R\pi_{*}{\mathbb{Q}}\cong\oplus R^{i}\pi_{*}{\mathbb{Q}}[-i]$, so the Leray spectral sequence corresponding to $\pi$ degenerates at $E_{2}$.
Corollary 3.5.
Let $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, and let $\pi\colon M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}$ be the projection from the universal degree $d$ $\operatorname{SL}$-character variety. Let $C_{g,n}:=2n^{2}(1-p_{n}^{-1})(g-1)$, where $p_{n}$ is the smallest prime factor of $n$. Then the truncation $\tau_{\leq C_{g,n}}R\pi_{*}{\mathbb{Q}}\cong\oplus_{i=0}^{C_{g,n}}R^{i}\pi_{*}{\mathbb{Q}}[-i]$. In particular, the Leray spectral sequence for $\pi$ degenerates at $E_{2}$ in degrees $\leq C_{g,n}$.
3.1. The moduli spaces
We now introduce the moduli spaces of Higgs bundles that appear on the Dolbeault side of the nonabelian Hodge correspondence.
Recall that a Higgs bundle on a curve $C$ is a pair $(E,\theta)$, where $E$ is a vector bundle on $C$ and $\theta\colon E\to E\otimes\omega_{C}$ a twisted endomorphism. The slope of a vector bundle $E$ is
$\mu(E):=\mathrm{degree}(E)/\mathrm{rank}(E).$
A Higgs bundle $(E,\theta)$ is stable (resp. semistable) if $\mu(F)<\mu(E)$ (resp. $\mu(F)\leq\mu(E))$ for every $\theta$-invariant subbundle $F\subset E$. When the degree and rank are coprime, a Higgs bundle is semistable if and only if it is stable. Let $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ denote the coarse moduli space of semistable Higgs bundles of degree $d$ and rank $n$ on $C$, as constructed by Simpson [49].
For moduli of $\operatorname{SL}$- and $\operatorname{PGL}$-Higgs bundles, we follow the definitions from [20, §2]. Fixing an auxiliary degree $d$ line bundle $L$ on $C$, the moduli space of $\operatorname{SL}$-Higgs bundles $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is the subspace of $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ parametrizing Higgs bundles $(E,\theta)$ such that $\det E\cong L$ and $\operatorname{tr}\theta=0$, when $\theta$ is viewed as an $\omega_{C}$-twisted endomorphism of $E$.
Definition 3.6.
The moduli space of $\operatorname{PGL}$-Higgs bundles is the geometric quotient
$M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)/\mathrm{Jac}(C)[n].$ (9)
Here, the group of $n$-torsion points $\mathrm{Jac}(C)[n]$ acts on $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ by $M\cdot(E,\theta):=(M\otimes E,\theta).$ For a different choice of degree $d$ line bundle $L^{\prime}$, there is a non-canonical isomorphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\cong M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L^{\prime})$ given by tensoring with an $n^{\mathrm{th}}$ root of $L^{-1}\otimes L^{\prime}\in\mathrm{Jac}(C)$. It follows that $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ is well-defined by (9) and independent of $L$.
These Higgs moduli spaces are related to the character varieties on the Betti side of the nonabelian Hodge correspondence, as follows [22, Theorem 4.3].
Fiberwise Nonabelian Hodge Correspondence.
Let $C$ be a smooth projective curve of genus $g$. For each $d$ and $n$ there is a homeomorphism $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{PGL}}_{B}(S_{g},n,d)$, and also
$M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)\cong M^{\operatorname{GL}}_{B}(S_{g,1},n,d),\quad\mbox{ and }\quad M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\cong M^{\operatorname{SL}}_{B}(S_{g,1},n,\deg L).$
3.2. Smoothness and purity for $d$ coprime to $n$
When $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, the Higgs moduli space $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ is a smooth algebraic variety [41, Proposition 7.4], and the corresponding fine moduli stack is a $\mathbb{C}^{*}$-gerbe over $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$. In this case, $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is also smooth.
Definition 3.7.
If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d):=[M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)/\mathrm{Jac}(C)[n]]$.
Thus, $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ is a smooth Deligne–Mumford stack with coarse space $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$.
The Higgs moduli space $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ is not proper, but it is proper over an affine space and shares some of the cohomological properties of a smooth and proper Deligne–Mumford stack, as we now recall.
The Hitchin morphism $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$ takes a Higgs bundle $(E,\theta)$ to the coefficients of the characteristic polynomial of $\theta$, viewed as a twisted endomorphism of $E$. The Hitchin morphism is proper [49, Theorem 6.11]. It is also equivariant with respect to the $\mathbb{G}_{m}$-action $t\cdot(E,\phi):=(E,t\phi)$ on the source and $t\cdot x\in H^{0}(C,\omega_{C}^{\otimes i}):=t^{i}x$ on the target. The Hitchin morphism is also invariant under the action of $\mathrm{Jac}(C)[n]$. It therefore factors through a proper morphism $p\colon M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\to\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. The $\mathbb{G}_{m}$-action likewise descends to an action on $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$ and, by the properness of $p$, any morphism $\mathbb{G}_{m}\to M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d),t\mapsto t\cdot x$ can be extended to a morphism from $\mathbb{A}^{1}$. The $\mathbb{G}_{m}$-fixed locus $M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d)^{\mathbb{G}_{m}}$ is proper, since it is a closed subspace of the fiber $p^{-1}(0)$.
Proposition 3.8.
If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then $H^{k}(\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d);\mathbb{Q})$ is pure of weight $k$.
Proof.
The proof is essentially identical to that of [18, Corollary 1.3.2]. ∎
Remark 3.9.
The character variety $M_{B}^{\operatorname{PGL}}(C,n,d)$ is homeomorphic to $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$, by the nonabelian Hodge correspondence, but this homeomorphism is not algebraic and the induced isomorphism on cohomology does not preserve weights. Indeed, the weight filtration on $H^{k}(M_{B}^{\operatorname{PGL}}(C,n,d);\mathbb{Q})$ is typically far from pure, and is the subject of the $P=W$ conjecture proposed in [5]. The $P=W$ conjecture for $\operatorname{PGL}_{n}$ follows from the corresponding conjecture for $\operatorname{GL}_{n}$, which is proved in [36, 16].
3.3. Universal Higgs moduli spaces
The definition of $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ readily generalizes to the relative setting, for a family of curves over a base. Also, there is an alternative characterization of the $\operatorname{PGL}$-Higgs moduli space as:
$\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\cong M_{\mathrm{Dol}}^{\operatorname{GL}}(C,n,d)/M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0).$ (10)
Here, $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)\cong\mathrm{Jac}(C)\times H^{0}(C,\omega_{C})$ is a group acting on $M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)$ by
$(M,s)\cdot(E,\theta)=(M\otimes E,s+\theta).$
Then (10) also generalizes to the relative setting. Thus, from the universal curve over $\mathcal{M}_{g}$, we obtain a universal $\operatorname{PGL}$-Higgs moduli space $M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$ By [49, Theorem 9.11 and Lemma 9.14], there is a homeomorphism from the relative $\operatorname{PGL}$-character variety to the relative $\operatorname{PGL}$-Higgs moduli space for an arbitrary family of curves. Simpson’s homeomorphism preserves the $C^{\infty}$-isomorphism class of the underlying $\operatorname{PGL}$-bundles, and hence maps degree $d$ character varieties to degree $d$ Higgs moduli spaces.
Universal $\operatorname{PGL}$ Nonabelian Hodge Correspondence.
There is a homeomorphism over $\mathcal{M}_{g}$
$M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\cong M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d).$
Corollary 3.10.
Let $\pi_{B}\colon\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ be the universal degree $d$ $\operatorname{PGL}$-character variety. Then the pushforward $R\pi_{B*}\mathbb{Q}$ splits as a sum of its cohomology sheaves if and only if the same is true for $\pi_{\mathrm{Dol}}\colon\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}.$
Proof.
We note that the pushforward of the constant sheaf $\mathbb{Q}$ from a Deligne–Mumford stack to its coarse space is the constant sheaf. Factoring the projection
$\pi_{B}\colon\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\xrightarrow{f_{B}}M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\xrightarrow{g_{B}}\mathcal{M}_{g},$
we see $R\pi_{B*}\mathbb{Q}\cong Rg_{B*}\mathbb{Q}$. Doing the same on the Dolbeault side shows $R\pi_{\mathrm{Dol}*}\mathbb{Q}=Rg_{\mathrm{Dol}*}\mathbb{Q}.$ The corollary follows, since $M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\cong M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ over $\mathcal{M}_{g}$. ∎
The universal Higgs moduli space $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ has a relative Hitchin morphism to the vector bundle whose fiber over $[C]$ is $\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})$. Furthermore, there is a natural relative compactification of the universal Higgs moduli space, whose construction is described in [8], expanding on [50, §11]. We briefly summarize its essential properties. Define a $\mathbb{G}_{m}$-action on $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\times\mathbb{A}^{1}$ by the standard scaling action on the first factor and $t\cdot x:=tx$ on the second. Let $M^{\circ}\subset M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ be the subset of points such that the $\mathbb{G}_{m}$-action does not have a limit at $\infty$. The quotient $M^{\circ}/\mathbb{G}_{m}]$ is a relative projective compactification of $M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)$ over $\mathcal{M}_{g}$. Moreover, both $M^{\circ}/\mathbb{G}_{m}]$ and the boundary divisor are smooth over $\mathcal{M}_{g}$.
Proposition 3.11.
If $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$ then there is a smooth relative compactification $\mathcal{X}$ of $\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)$ over $\mathcal{M}_{g}$ whose boundary $\mathcal{X}\smallsetminus\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)$ is a smooth relative divisor.
Remark 3.12.
The argument proving the degeneration of the Leray spectral sequence for $\mathcal{M}_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ applies equally to $M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}$. Indeed, the $\operatorname{GL}$ spaces satisfy the natural analogs of Propositions 3.8 and 3.11 for essentially the same reasons.
3.4. Proof of Theorem 3.1 using Saito’s theory of Hodge modules
To handle constructible sheaves equipped with Hodge structures, we use Saito’s theory of mixed Hodge modules [45]. The key features are as follows. There is a category of mixed Hodge modules $\operatorname{MHM}(X)$ on each complex variety $X$, with a standard six-functor formalism. There is a functor
$\mathrm{rat}\colon D^{b}\operatorname{MHM}(X)\to D^{b}_{c}(\mathbb{Q}_{X})$ (11)
taking a mixed Hodge module to its underlying rational structure, which is compatible with the six functors [45, Theorem 0.1]. Moreover, $\operatorname{MHM}(\mathrm{pt})$ is the category of polarizable $\mathbb{Q}$-mixed Hodge structures [45, (4.2.12)], and pushing forward the trivial Hodge module $\mathbb{Q}^{H}_{X}$ to a point we recover Deligne’s mixed Hodge structure on $H^{*}(X;\mathbb{Q})$ [45, p. 328].
Remark 3.13.
Deligne’s theory of mixed Hodge structures extends to Deligne–Mumford stacks and, more generally, simplicial objects in the category of algebraic varieties [11]. Saito’s theory likewise extends to Deligne–Mumford stacks, as explained in [51]. We use the basic properties of mixed Hodge structures and mixed Hodge modules on Deligne–Mumford stacks without further comment.
We also note that all of the Deligne–Mumford stacks we consider are global quotients of algebraic varieties by finite groups. Thus, our arguments can be rephrased in terms of equivariant mixed Hodge modules and invariant subspaces of mixed Hodge structures on varieties with finite group actions without any other essential changes.
The needed input from the theory of mixed Hodge modules is the following lemma:
Lemma 3.14.
If $\pi\colon U\to S$ is a smooth morphism of smooth complex Deligne–Mumford stacks such that the cohomology of each fiber $H^{k}(U_{s};\mathbb{Q})$ has a pure Hodge structure of weight $k$ for each degree $k$ and $s\in S$, then $R\pi_{*}\mathbb{Q}_{U}\cong\oplus_{i}R^{i}\pi_{*}\mathbb{Q}_{U}[-i]$. In particular, the corresponding Leray spectral sequence degenerates at $E_{2}$.
Proof.
Our goal is to show that the (derived) pushforward $\pi_{*}\mathbb{Q}^{H}_{U}\in D^{b}\operatorname{MHM}(S)$ is pure of weight 0 [45, §4.5], i.e., that $\mathrm{Gr}^{W}_{i}\mathcal{H}^{j}(\pi_{*}\mathbb{Q}^{H}_{U})=0$ for $i\neq j$. This condition guarantees that $\pi_{*}\mathbb{Q}^{H}_{U}\cong\oplus_{i}\mathcal{H}^{i}(\pi_{*}\mathbb{Q}^{H}_{U})[-i]\in D^{b}\operatorname{MHM}(S)$ [45, (4.5.4)], and then applying the rat functor (11), one recovers the same identity on the level of constructible sheaves.
To prove the purity condition, it suffices to check that for every $s\in S$ and $k\in\mathbb{Z}$ we have that $\mathcal{H}^{k}i_{s}^{*}\pi_{*}\mathbb{Q}^{H}_{U}$ is of weight $\leq k$ and $\mathcal{H}^{k}i_{s}^{!}\pi_{*}\mathbb{Q}_{U}^{H}$ is of weight $\geq k$ [45, (4.6.1)]. We start with the identity [45, (4.4.3)]
$i_{s}^{!}\pi_{*}\mathbb{Q}^{H}_{U}\cong\pi_{U_{s}*}i^{!}_{U_{s}}\mathbb{Q}_{U}^{H}.$
For the first assertion, for $d:=\dim S$, note that $i_{s}^{*}=i_{s}^{!}(-d)[-2d]$ by smoothness of $S$ and $i_{U_{s}}^{*}=i_{U_{s}}^{!}(-d)[-2d]$ by smoothness of $S$ and $f$ [45, (4.4.1)], and thus
$i_{s}^{*}\pi_{*}\mathbb{Q}^{H}_{U}=i_{s}^{!}\pi_{*}\mathbb{Q}^{H}_{U}(-d)[-2d]=\pi_{U_{s}*}i^{!}_{U_{s}}\mathbb{Q}^{H}_{U}(-d)[-2d]=\pi_{U_{s}*}i^{*}_{U_{s}}\mathbb{Q}^{H}_{U}=H^{*}(U_{s};\mathbb{Q}),$
which is of weight 0 by assumption. Similarly, for the second assertion,
$i_{s}^{!}\pi_{*}\mathbb{Q}^{H}_{U}=\pi_{U_{s}*}i^{!}_{U_{s}}\mathbb{Q}^{H}_{U}=\pi_{U_{s}*}i_{U_{s}}^{*}\mathbb{Q}^{H}_{U}(-d)[-2d]=H^{*-2d}(U_{s};\mathbb{Q})(-d),$
which is again of weight 0 by assumption. ∎
First proof of Theorem 3.1.
By Corollary 3.10 it suffices to prove the statement on the Dolbeault side, for $\pi\colon\mathcal{M}_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$. Recall that $\pi$ is a smooth morphism [41, Proposition 7.4] whose fibers have pure cohomology, by Proposition 3.8. The theorem then follows from Lemma 3.14.
3.5. Proof of Theorem 3.1 using classical Hodge theory and derived categories
In this subsection, we give a somewhat longer but more elementary proof of Theorem 3.1. The strategy is the same, except that Saito’s general result on the splitting of the pushforward of a pure Hodge module (Lemma 3.14) is replaced by classical Hodge theory on the fibers and a general observation on splitting in triangulated categories (Lemma 3.15).
Second proof of Theorem 3.1.
Let $U:=\mathcal{M}^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)$. By Corollary 3.10 it suffices to prove the result for $\pi\colon U\to\mathcal{M}_{g}$. By Proposition 3.11, there is a smooth relative compactification $X=U\sqcup Z$, where $Z$ is a divisor in $X$ that is also smooth over $\mathcal{M}_{g}$.
Then there is an exact triangle
$i_{*}i^{!}\mathbb{Q}_{X}\to\mathbb{Q}_{X}\to Rj_{*}j^{-1}\mathbb{Q}_{X}\xrightarrow{+1}$
in $D^{b}_{c}(\mathbb{Q}_{X})$ [13, p. 45], where $j^{-1}\mathbb{Q}_{X}=\mathbb{Q}_{U}$. Since $X$ and $Z$ are smooth, and $Z$ is of complex codimension 1, we have $i^{!}\mathbb{Q}_{X}=\mathbb{Q}_{Z}[-2]$ [13, Proposition 3.2.11]. Pushing forward along the map $\pi\colon X\to\mathcal{M}_{g}$, we obtain an exact triangle
$R\pi_{Z*}\mathbb{Q}_{Z}[-2]\to R\pi_{X*}\mathbb{Q}_{X}\to R\pi_{U*}\mathbb{Q}_{U}\xrightarrow{+1}$ (12)
whose long exact sequence of cohomology is the relative Gysin sequence. Restricting to the fiber over any point $s\in\mathcal{M}_{g}$, using proper base change [13, Theorem 2.3.26] to identify $(R\pi_{X*}\mathbb{Q}_{X})_{s}\cong H^{*}(X_{s},\mathbb{Q})$ and similarly for $Z$, and taking the long exact sequence of cohomology, we obtain the Gysin sequence for the smooth divisor $Z_{s}\subset X_{s}$
$\cdots\rightarrow H^{i-2}(Z_{s};\mathbb{Q})(-1)\to H^{i}(X_{s};\mathbb{Q})\to H^{i}(U_{s};\mathbb{Q})\rightarrow\cdots.$
This is an exact sequence of mixed Hodge structures [11, (9.2.1.2)]. By Proposition 3.8, $H^{i}(U_{s})$ is pure of weight $i$. The same is true for the cohomology groups of $X_{s}$ and $Z_{s}$, since both of these spaces are smooth and proper. In particular, this implies that the long exact sequence above breaks into short exact sequences. Now, since this is true on each fiber and vanishing of a map of sheaves can be checked on fibers, we conclude that the relative Gysin sequence splits into short exact sequences
$0\to R^{i-2}\pi_{Z*}\mathbb{Q}_{Z}\to R^{i}\pi_{X*}\mathbb{Q}_{X}\to R^{i}\pi_{U*}\mathbb{Q}_{U}\to 0.$
Since $\pi_{X}\colon X\to\mathcal{M}_{g}$ is a smooth and projective map to a smooth base, we have
$R\pi_{*}\mathbb{Q}_{X}\cong\oplus R^{i}\pi_{*}\mathbb{Q}_{X}[-i].$
Note that $R^{i}\pi_{*}\mathbb{Q}_{X}$ is a semisimple local system, by [9, Proposition 2.1] and [10, Cor. 4.2.9(a)]. The same being true for $\pi_{Z}$, we conclude that each $R^{i}\pi_{U*}\mathbb{Q}_{U}$ is also a local system, and each short exact sequence above is split, by semisimplicity of the middle factor.
Theorem 3.1 therefore follows from Lemma 3.15 below, applied to the standard $t$-structure on $D^{b}_{c}(\mathbb{Q}_{\mathcal{M}_{g}})$ and the triangle in equation (12) (cf. [13, Example 5.1.3]). ∎
Lemma 3.15.
Let $\mathcal{D}$ be a triangulated category with a bounded $t$-structure. Say $F\in\mathcal{D}$ is split if $F\cong\oplus\mathcal{H}^{i}(F)[-i]$. Given a distinguished triangle $A\to B\to C\xrightarrow{+1}$ such that $B$ is split, $C\to A[1]$ induces a trivial map on cohomology, and the resulting short exact sequences
$0\to\mathcal{H}^{i}(A)\to\mathcal{H}^{i}(B)\to\mathcal{H}^{i}(C)\to 0$
are split, then $C$ is also split.
Proof.
Choose the maximum integer $m$ such that at least one of $A,B,C$ has a nontrivial $t$-cohomology group $\mathcal{H}^{m}$; this maximum exists by assumption that the $t$-structure is bounded, and since these splitting properties are clearly compatible with shift functors, we may assume without loss of generality that $m=0$. Then our first goal is to show that $C\cong C_{<0}\oplus\mathcal{H}^{0}(C)$, where $C_{<0}:=\tau^{\leq-1}C$ is the truncation in the definition of the $t$-structure. In other words, there is an exact triangle
$C_{<0}\to C\to C_{\geq 0}\xrightarrow{+1}$
where the third term is $\mathcal{H}^{0}(C)$ by assumption that $C$ has trivial cohomology in degrees >0>0, and we aim to prove that the connecting morphism $C_{\geq 0}=\mathcal{H}^{0}(C)\rightarrow C_{<0}[1]$ is trivial.
We consider the diagram
$\mathcal{H}^{0}(B),$$B_{<0}[1]$$\mathcal{H}^{0}(C)$$C_{<0}[1]$$\mathcal{H}^{0}(A)[1]$$\scriptstyle{0}$$\scriptstyle{g}$$\scriptstyle{f}$$\scriptstyle{0}$
where the square is induced by applying to $B\to C$ the natural transformation $\tau^{\geq 0}\to\tau^{<0}[1]$ from the truncation exact triangle, and the left column comes from a rotation of the triangle coming from the short exact sequence
$0\to\mathcal{H}^{0}(A)\to\mathcal{H}^{0}(B)\to\mathcal{H}^{0}(C)\to 0$
that is split by assumption, making the bottom arrow 0. Similarly, the top map is trivial by assumption that $B$ is split. Considering the exact sequence
$\mathrm{Hom}(\mathcal{H}^{0}(A)[1],C_{<0}[1])\xrightarrow{\circ 0}\mathrm{Hom}(\mathcal{H}^{0}(C),C_{<0}[1])\xrightarrow{\circ g}\mathrm{Hom}(\mathcal{H}^{0}(B),C_{<0}[1]),$
the second map must be injective; then since commutativity of the square shows that $f\circ g=0$, we must have $f=0$, as needed.
We now claim that there is a triangle
$A_{<0}\to B_{<0}\to C_{<0}\xrightarrow{+1}$
satisfying the hypotheses of the lemma. To prove this, we apply the 9-lemma [2, Proposition 1.1.11] to the commutative square coming from applying $id\to\tau^{\geq 0}$ to $A\to B$:
[…]
This gives a $3\times 3$ diagram all of whose rows and columns are distinguished triangles:
[…]
(The fact that the second column forms a short exact sequence uses our assumption that the long exact sequence of cohomology breaks into short exact sequences.) This gives a rotated form of our desired triangle. Furthermore, $B_{<0}$ is split, as
$\tau_{<0}(\oplus_{i\leq 0}\mathcal{H}^{i}(B)[-i])=\oplus_{i<0}\mathcal{H}^{i}(B)[-i],$
and the long exact sequence of cohomology for this triangle is that of the original triangle in degree <0<0 and 0 otherwise, thus satisfying the conditions that $C_{<0}\to A_{<0}[1]$ induces a trivial map on cohomology and the resulting short exact sequences are split.
Thus we can continue to apply the argument to split $C_{<0}\cong\mathcal{H}^{-1}(C)[1]\oplus C_{<-1}$, etc. Again using the assumption that the $t$-structure is bounded, there is some integer $n$ such that $C$ has no cohomology in degrees <n<n, thus $C_{<n}=0$, and so the inductively constructed isomorphism $C\cong\oplus_{i\geq n}\mathcal{H}^{i}(C)[-i]\oplus C_{<n}$ in fact gives a splitting of $C$. ∎
3.6. Proof of Corollaries 3.4 and 3.5: the case of $\operatorname{GL}$ and $\operatorname{SL}$
Here, we deduce the splitting results for the $\operatorname{GL}$ and $\operatorname{SL}$ relative character varieties from the result for $\operatorname{PGL}$ (Theorem 3.1), using the morphisms described in Proposition 2.5. For $\operatorname{SL}$, this is more or less immediate from the facts that the $\operatorname{PGL}$ space is a finite group quotient of the $\operatorname{SL}$ space and the group action is trivial on low degree cohomology. We deduce the result for $\operatorname{GL}$ from the $\operatorname{SL}$ result and a relative Künneth formula (Lemma 3.16).
Proof of Corollary 3.5.
First, recall that the splitting in Theorem 3.1 is stated for the universal $\operatorname{PGL}$-character variety over $\mathcal{M}_{g}$. By smooth base change [13, Theorem 3.2.13(ii)], the analogous splitting holds for any family of curves induced by a smooth morphism to $\mathcal{M}_{g}$. In particular, this holds for the universal curve over $\mathcal{M}_{g,1}$.
Let $\Gamma:=\mathrm{Hom}(\pi_{1}(S_{g},1),\mu_{n})$. By Proposition 2.5, the relative $\operatorname{PGL}$-character variety $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g,1},n,d)$ over $\mathcal{M}_{g,1}$ is the quotient of $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)$ by the relative group object $\widetilde{\Gamma}:=\mathcal{T}_{g,1}\times_{\operatorname{Mod}_{g,1}}\Gamma$. In particular, taking cohomology of a fiber, we have
$H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})\cong H^{*}(M_{B}^{\operatorname{SL}}(S_{g,1},n,d);\mathbb{Q})^{\Gamma}.$
It follows that $R^{i}\pi_{\operatorname{PGL}*}\mathbb{Q}\to(R^{i}\pi_{\operatorname{SL}*}\mathbb{Q})^{\widetilde{\Gamma}}$ is an isomorphism for all $i$, and
$R\pi_{\operatorname{PGL}*}\mathbb{Q}\cong(R\pi_{\operatorname{SL}*}\mathbb{Q})^{\widetilde{\Gamma}}$. (13)
Let $C_{g,n}:=2n^{2}(1-p_{n}^{-1})(g-1)$, where $p_{n}$ is the smallest prime factor of $n$. We claim that $R^{i}\pi_{\operatorname{SL}*}\mathbb{Q}$ is $\widetilde{\Gamma}-$invariant for $i\leq C_{g,n}$. The corollary then follows from the claim, combined with (13) and Theorem 3.1.
It remains to prove the claim. By [17, Theorem 1.4], we have
$H^{i}(M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d);\mathbb{Q})\cong H^{i}(M_{\mathrm{Dol}}^{\operatorname{SL}}(C,n,d);\mathbb{Q})\mbox{ for }i\leq C_{g,n}.$ (14)
By the nonabelian Hodge correspondence for $\operatorname{SL}$ and $\operatorname{PGL}$, we have homeomorphisms
$M_{B}^{\operatorname{PGL}}(S_{g},n,d)\cong M_{\mathrm{Dol}}^{\operatorname{PGL}}(C,n,d)\quad\mbox{ and }\quad M_{B}^{\operatorname{SL}}(S_{g,1},n,d)\cong M_{\mathrm{Dol}}^{\operatorname{SL}}(C,n,d).$ (15)
Combining (14) and (15), we conclude that $H^{i}(M_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})\cong H^{i}(M_{B}^{\operatorname{SL}}(S_{g,1},n,d);\mathbb{Q})$, and $R^{i}\pi_{\operatorname{PGL}*}\mathbb{Q}\xrightarrow{\sim}(R^{i}\pi_{\operatorname{SL}*}\mathbb{Q})$ and $(R^{i}\pi_{\operatorname{SL}*}\mathbb{Q})$ is $\widetilde{\Gamma}-$invariant, for $i\leq C_{g,n}$, as claimed. ∎
In the proof of Corollary 3.4, we use the following lemma.
Lemma 3.16.
Let $\pi_{i}\colon X_{i}\to S$, $i=1,2$, be a submersion of complex manifolds, with $\pi\colon X_{1}\times_{S}X_{2}\to S$. Suppose that each $R^{j}\pi_{2*}\mathbb{Q}$ is locally trivial. Then
$R\pi_{*}\mathbb{Q}\cong R\pi_{1*}\mathbb{Q}\otimes R\pi_{2*}\mathbb{Q}$
in $D^{b}_{c}(\mathbb{Q}_{S})$. The same holds if the base $S$ is a smooth complex Deligne–Mumford stack.
Proof.
All of the functors in this proof are derived; for simplicity we omit this from the notation. Also, in this proof (only) we use $\dim$ to denote $\dim_{\mathbb{R}}$. We start with the case that $S$ is a manifold. The strategy is to apply Verdier duality to the proper relative Künneth formula [25, Exercise II.18(i)]
$\pi_{!}(\mathbb{Q})=\pi_{1!}\mathbb{Q}\otimes\pi_{2!}\mathbb{Q}.$ (16)
On the left side we have
$D\pi_{!}\mathbb{Q}=\pi_{*}D\mathbb{Q}=\pi_{*}\mathbb{Q}[\dim X_{1}\times_{S}X_{2}]$
[13, Theorem 3.2.17(iii) and Proposition 3.3.7]. On the right side we have
$D(\pi_{1!}\mathbb{Q}\otimes\pi_{2!}\mathbb{Q})=\mathcal{H}om(\pi_{1!}\mathbb{Q},\mathcal{H}om(\pi_{2!}\mathbb{Q},\omega_{S}))=\mathcal{H}om(\pi_{1!}\mathbb{Q},\pi_{2*}D\mathbb{Q}),$
where for the first equality we use [25, Proposition 2.6.3(ii)]. Identifying $D\mathbb{Q}=\mathbb{Q}[\dim X_{2}]$ and using the assumption that the sheaves $R^{j}\pi_{2*}\mathbb{Q}$ are locally trivial we see from [7, p. 624] that
$D(\pi_{1!}\mathbb{Q}\otimes\pi_{2!}\mathbb{Q})=\mathcal{H}om(\pi_{1!}\mathbb{Q},\mathbb{Q}[\dim S])\otimes\pi_{2*}\mathbb{Q}[\dim X_{2}-\dim S]$
$D\pi_{1!}\mathbb{Q}\otimes\pi_{2*}\mathbb{Q}[\dim X_{2}-\dim S]=(\pi_{1*}\mathbb{Q}\otimes\pi_{2*}\mathbb{Q})[\dim X_{1}+\dim X_{2}-\dim S].$
Thus the needed identity is obtained by applying Verdier duality to (16) and shifting by $\dim X_{1}\times_{S}X_{2}$.
If $S$ is a smooth Deligne–Mumford stack, we choose an étale cover $p:S^{\prime}\to S$ by a smooth variety $S^{\prime}$. By smooth base change, the canonical morphism $R\pi_{1*}\mathbb{Q}\otimes R\pi_{2*}\mathbb{Q}\to R\pi_{*}\mathbb{Q}$ becomes an isomorphism after pulling back to $S^{\prime}$. In particular, the mapping cone $C$ satisfies $p^{*}C=0$, and thus $C=0$. ∎
Proof of Corollary 3.4.
By Proposition 2.5, the character variety $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d)$ is a quotient of $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\times M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)$ by the relative group object $\widetilde{\Gamma}$ with fiber $\Gamma=\mathrm{Hom}(\pi_{1}(S_{g,1}),\mu_{n})$ over $\mathcal{M}_{g,1}$. We have an isomorphism $R\pi_{\operatorname{GL}*}\mathbb{Q}=(R\pi_{\operatorname{GL}_{1}\times\operatorname{SL}}\mathbb{Q})^{\widetilde{\Gamma}}$, as in (13). By Lemma 3.16, using that $M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)$ and $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$ are locally trivial fibrations with smooth fibers, we see that
$R\pi_{\operatorname{GL}_{1}\times\operatorname{SL}}\mathbb{Q}=R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}\otimes R\pi_{\operatorname{SL}*}\mathbb{Q}.$
To show the splitting of $R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$, we use Simpson’s ($d=0$) nonabelian Hodge theorem $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\cong M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$ [49, Proposition 6.8 and Theorem 7.18] and the fact that, as noted in Remark 3.12, the proof of Theorem 3.1 holds for $M_{\mathrm{Dol}}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)$. Furthermore, $\Gamma$ acts trivially on the cohomology of the fibers. This is because the action on the fibers can be extended to the multiplication action of the connected group $\mathrm{Hom}(\pi_{1}(S_{g,1}),\mathbb{C}^{*})\cong(\mathbb{C}^{*})^{2g}$ on itself. Then, since $M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},1,0)\to\mathcal{M}_{g,1}$ is topologically locally trivial, we have $(R\pi_{\operatorname{GL}_{1}*}\mathbb{Q})^{\widetilde{\Gamma}}=R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$. Thus, we conclude that
$R\pi_{\operatorname{GL}*}\mathbb{Q}=(R\pi_{\operatorname{GL}_{1}\times\operatorname{SL}}\mathbb{Q})^{\widetilde{\Gamma}}=(R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}\otimes R\pi_{\operatorname{SL}*}\mathbb{Q})^{\widetilde{\Gamma}}=R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}\otimes(R\pi_{\operatorname{SL}*}\mathbb{Q})^{\widetilde{\Gamma}}$. (17)
Applying the splitting for $R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$ above and the splitting for $(R\pi_{\operatorname{SL}*}\mathbb{Q})^{\widetilde{\Gamma}}\cong R\pi_{\operatorname{PGL}*}\mathbb{Q}$ as described in Corollary 3.5, we get the desired splitting for $R\pi_{\operatorname{GL}*}\mathbb{Q}$. ∎
4. Computations of the stable rational cohomology
In this section, we prove Theorem 1.1 and Theorem 1.4 by combining the degeneration of the Leray spectral sequence at $E_{2}$ (Theorem 3.1 and Corollaries 3.43.5) with previously known results on the cohomology of the fibers and the cohomology of $\mathcal{M}_{g}$ and $\mathcal{M}_{g,1}$ with coefficients in symmetric and exterior powers of the symplectic local systems.
4.1. Monodromy representations
The structure map of the universal $\operatorname{PGL}$-character variety $\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ is a topologically locally trivial fiber bundle, as are
$M_{B}^{\operatorname{SL}}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}\quad\mbox{ and }\quad M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}.$
We now study the local systems on $\mathcal{M}_{g}$ and $\mathcal{M}_{g,1}$ that arise from the cohomology of the fibers with their respective monodromy representations.
As in §2, we write $X_{g,n,d}$ for the subspace of $\mathrm{Hom}(\pi_{1}(S_{g}),\operatorname{PGL}_{n}(\mathbb{C}))$ parametrizing maps whose obstruction to lifting to $\operatorname{SL}_{n}(\mathbb{C})$ is $d\in\mathbb{Z}/n\mathbb{Z}$. Let $\widetilde{S}_{g}$ be the universal cover of $S_{g}$.
Definition 4.1.
The tautological $\operatorname{PGL}_{n}$-bundle over $S_{g}\times\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d)$ is
$\mathcal{P}:=\big(\widetilde{S}_{g}\times X_{g,n,d}\times\operatorname{PGL}_{n}(\mathbb{C})\big)/\big(\pi_{1}(S_{g},p)\times\operatorname{PGL}_{n}(\mathbb{C})\big).$
Here, the action of $\pi_{1}(S_{g},p)\times\operatorname{PGL}_{n}(\mathbb{C})$ can be described as
$(\gamma,M)\cdot(x,\phi,N):=(\gamma x,M\phi M^{-1},M\phi(\gamma)N).$
The structure map to $S_{g}\times\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ is induced by projection to the first two factors.
Remark 4.2.
The universal $\operatorname{PGL}_{n}$ bundle on $S_{g}\times M_{B}^{\operatorname{GL}}(S_{g,1},n,d)$ described in [21, §1] is the pullback of $\mathcal{P}$ by the quotient map $M_{B}^{\operatorname{GL}}(S_{g,1},n,d)\to\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$, as is the universal $\operatorname{PGL}_{n}$-bundle on $S_{g}\times M_{B}^{\operatorname{SL}}(S_{g,1},n,d)$ in [19, §4.1].
The generators for $H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})$ are the Künneth components of the Chern classes of this tautological bundle. We make the following (basis-free) definition:
Definition 4.3.
Let $p_{k}$ be the projection of $S_{g}\times\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ onto the $k^{\textrm{th}}$ component. For $i\in\{0,1,2\}$ and $j\in\mathbb{Z}_{\geq 2}$ we define the Künneth component vector space
$V_{i,j}:=\{p_{2*}(c_{j}(\mathcal{P})\cup p_{1}^{*}(x)):x\in H^{i}(S_{g};\mathbb{Q})\}\subset H^{i+2j-2}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q}).$
Proposition 4.4.
The subspaces $V_{i,j}$ for $0\leq i\leq 2$ and $2\leq j\leq n$ generate the ring $H^{*}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Q})$ with no relations in degrees less than or equal to 2(g1)(n1)+22(g-1)(n-1)+2, and $\dim V_{i,j}=\dim H^{i}(S_{g};\mathbb{Q})$.
Proof.
The generation statement is a consequence of Markman’s tautological generation result for $M_{\mathrm{Dol}}^{\operatorname{GL}}(C,n,d)$ [32, Theorem 7], as reinterpreted on the Betti side by Hausel and Thaddeus [21, (5.1)]. Picking a homogeneous basis of $H^{*}(S_{g};\mathbb{Q})$, there are no relations in degrees $\leq 2(g-1)(n-1)+2$ among the images of these elements in $V_{*,j}$ by [19, Lemma 4.1.12]. ∎
Remark 4.5.
In later work [33, Theorem 3], Markman strengthened the generation result from rational to integral cohomology. In particular, in degrees $\leq 2(g-1)(n-1)+2$ one concludes that $H^{*}(M^{\operatorname{GL}}_{B}(S_{g,1},n,d);\mathbb{Z})$ is torsion-free. Then using the quotient description of Proposition 2.5, we conclude that the cohomology groups of the fibers $H^{*}(M^{\operatorname{SL}}_{B}(S_{g,1},n,d);\mathbb{Z}[n^{-1}])$ and $H^{*}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Z}[n^{-1}])$ are torsion-free in the same degree range. However, in order to deduce a result about integral cohomology of the universal spaces, one would also need torsion-freeness results for the various cohomology groups of $\mathcal{M}_{g}$ and $\mathcal{M}_{g,1}$ with symplectic coefficients.
Proposition 4.6.
The Chern classes $c_{i}(\mathcal{P})\in H^{2i}(S_{g}\times\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})$ of the universal $\operatorname{PGL}_{n}$-bundle $\mathcal{P}$ on $S_{g}\times\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)$ defined in Definition 4.1 are fixed by the mapping class group action.
Proof.
We follow the strategy of [21, (5.1)]. For any element of $\operatorname{Mod}_{g}$, we may choose a representative diffeomorphism $f$ of $S_{g}$ fixing a chosen point $p$. It is then straightforward to check that $\mathcal{P}$ and $f^{*}\mathcal{P}$ are both $\operatorname{PGL}_{n}$-equivariant flat $\operatorname{PGL}_{n}$-bundles over $S_{g}\times X_{g,n,d}$ trivialized over $\{p\}\times X_{g,n,d}$ such that the monodromy representation of the restriction to $S_{g}\times\{\phi\}$ is given by $\phi$, and thus $\mathcal{P}\cong f^{*}\mathcal{P}$.
Remark 4.7.
A cleaner but more technical proof of the previous proposition goes as follows: take a simplicial set with geometric realization homotopic to $S_{g}$, and view this as a simplicial scheme $X$ by assigning $\mathrm{Spec}\,\mathbb{C}$ to each element. Then
$\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d)\cong\underline{\mathrm{Hom}}(X,\mathrm{B}\operatorname{PGL}_{n}),$
where $\underline{\mathrm{Hom}}$ is the internal Hom in the category of simplicial schemes. From this perspective, it is clear that classes pulled back under the evaluation map
$X\times{\underline{\mathrm{Hom}}}(X,\mathrm{B}\operatorname{PGL}_{n})\to\mathrm{B}\operatorname{PGL}_{n}$
are canonical, and in particular invariant under homeomorphisms of $S_{g}$. (See [47] for more discussion.)
Corollary 4.8.
The subspaces $V_{0,j},V_{2,j}\subset H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})$ are fixed by $\operatorname{Mod}_{g}$, and $V_{1,j}$ is the standard symplectic representation.
Proof.
By construction, we have a $\operatorname{Mod}_{g}$-fixed vector
$c_{j}(\mathcal{P})\in\bigoplus_{i=0}^{2}H^{i}(S_{g};\mathbb{Q})\otimes V_{2-i,j}\subset H^{*}(S_{g};\mathbb{Q})\otimes H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q}).$
The projection of $c_{j}(\mathcal{P})$ in each $H^{i}(S_{g};\mathbb{Q})\otimes V_{2-i,j}$ is then also fixed (and nonzero by Proposition 4.4). Since $\operatorname{Mod}_{g}$ acts irreducibly on each $H^{i}(S_{g};\mathbb{Q})$ and the dimensions match (again by Proposition 4.4), we conclude that $V_{2-i,j}=H^{i}(S_{g};\mathbb{Q})^{\vee}$. In other words, $V_{0,j}$ and $V_{2,j}$ are the trivial representation, and $V_{1,j}$ the standard symplectic representation. ∎
The previous corollary, combined with Proposition 4.4, suffices to describe all the local systems $R^{i}\pi_{*}\mathbb{Q}$ for $\pi\colon\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$. Then the $\operatorname{PGL}$ result, together with the proof of Corollary 3.5, determines the local systems appearing for $\operatorname{SL}$. The case of $\operatorname{GL}$ follows from the case of $\operatorname{SL}$, the proof of Corollary 3.4, and one last observation:
Proposition 4.9.
The $\operatorname{Mod}_{g}$ representation $H^{i}(M_{B}^{\operatorname{GL}}(S_{g},1,0);\mathbb{Q})$ is the $i^{\textrm{th}}$ exterior power of the standard symplectic representation.
Proof.
We note that
$M_{B}(S_{g},1,0)=\mathrm{Hom}(\pi_{1}(S_{g}),\mathbb{C}^{*})=\mathrm{Hom}(H_{1}(S_{g};\mathbb{Z});\mathbb{C}^{*})\cong(\mathbb{C}^{*})^{2g}$
with $\operatorname{Mod}_{g,1}$ acting through the symplectic representation on $H_{1}(S_{g};\mathbb{Z}).$
4.2. Stable rational cohomology of $\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)$
We relate $H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d);\mathbb{Q})$ to the cohomology of $\mathcal{M}_{g}$ with coefficients in exterior powers of direct sums of copies of the symplectic local system. To simplify the notation of the relevant local systems on $\mathcal{M}_{g}$, recall the conventions (2) from the introduction:
$V_{\circ}:=\mathbb{Q}\oplus\mathbb{Q}[-2],\quad\quad W_{n}:=\mathbb{Q}\oplus\mathbb{Q}[-2]\oplus\cdots\oplus\mathbb{Q}[2-2n],\quad\quad\quad\widetilde{W}_{n}:=W_{n}/\mathbb{Q}.$ (18)
Proof of Theorem 1.1.
From Theorem 1.3 we have
$H^{k}(M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d);\mathbb{Q})\cong\bigoplus\limits_{p+q=k}H^{p}(\mathcal{M}_{g};H^{q}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Q})).$ (19)
By Proposition 4.4 we have
$H^{*}(\mathcal{M}^{\operatorname{PGL}}_{B}(S_{g},n,d);\mathbb{Q})\cong\wedge^{\bullet}(\oplus_{j=2}^{n}V_{1,j})\otimes\operatorname{Sym}^{\bullet}(\oplus_{j=2}^{n}V_{0,j}\oplus V_{2,j})$
in degrees $\leq 2(g-1)(n-1)+2$. Using the notation (18) and the description of the $\operatorname{Mod}_{g}$ action in Corollary 4.8, we can rewrite the subspaces $V_{i,j}$ of $H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(S_{g},n,d);\mathbb{Q})$ as the following graded modules, where $V_{g}=H^{1}(S_{g};\mathbb{Q})$, as in (1), in degree 1:
$\oplus_{j=2}^{n}V_{1,j}=V_{g}\otimes\widetilde{W}_{n},\quad\oplus_{j=2}^{n}V_{0,j}\oplus V_{2,j}=V_{\circ}\otimes\widetilde{W}_{n}.$
Hence we conclude that in this range
$H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d);\mathbb{Q})=H^{*}\big(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n})\big)\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n}).\qed$
Proof of Corollary 1.2.
We write $S_{\langle\lambda\rangle}(V_{g})$ for the irreducible representation of $\operatorname{Sp}(2g,\mathbb{Q})$ corresponding to the partition $\lambda$ (in degree 0). We begin by partially decomposing
$\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n})=\bigotimes_{i=1}^{n-1}\wedge^{\bullet}(V_{g}[-2i]),$
where
[…]
Then in degree $\leq 3g$ the cohomology $H^{*}\big(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n})\big)$ splits (independently of $g$) as a direct sum of terms of the form
$H^{*}(\mathcal{M}_{g};S_{\langle 1^{r_{1}}\rangle}(V_{g})[s_{1}]\otimes\cdots\otimes S_{\langle 1^{r_{n-1}}\rangle}(V_{g})[s_{n-1}])$ (21)
with $s_{i}=(1+2i)r_{i}^{\prime}$ for some $r_{i}^{\prime}\geq r_{i}$ satisfying $r_{i}^{\prime}\equiv r_{i}$ mod $2$. Such a term has highest weight $\sum\langle 1^{r_{i}}\rangle$, and so each irreducible piece $S_{\langle\lambda\rangle}(V_{g})$ satisfies $|\lambda|\leq\sum r_{i}$. Furthermore, using the Littlewood-Richardson rule to compute the decomposition into irreducibles (see [28]), one sees that the set of irreducibles appearing in the decomposition of each term (21) in degrees $k\leq 3g$ is independent of $g$; also each $S_{\langle\lambda\rangle}(V_{g})$ appearing satisfies $|\lambda|\equiv\sum r_{i}$ mod $2$. Now, Looijenga [30, Theorem 1.1] has calculated the stable cohomology of $H^{*}(\mathcal{M}_{g};S_{\langle\lambda\rangle}(V_{g}))$ in degree $\leq(2g-2)/3-|\lambda|$. Since $\sum s_{i}\geq\sum r_{i}\geq|\lambda|$, we note that the contribution of each irreducible piece of each term (21) to $H^{k}(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))$ comes from Looijenga’s stable range when $k\leq(2g-2)/3$. In particular, this is independent of $g$.
Finally, when $H^{k}(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))$ is decomposed as a sum of terms
$H^{k-\sum s_{i}}(\mathcal{M}_{g};S_{\langle\lambda\rangle}(V_{g}))$
with $|\lambda|\equiv\sum r_{i}\equiv\sum s_{i}$ mod $2$, it follows from Looijenga’s description that each such term is 0 when $k$ is odd. The same then holds for the odd cohomology of
$H^{*}(\mathcal{M}_{B}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d);\mathbb{Q})=H^{*}(\mathcal{M}_{g},\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n}))\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n})$
since the second term is nontrivial only in even degree.
Finally, since $\widetilde{W}_{n}$ and $\widetilde{W}_{n+1}$ agree up to degree $2n$, we note that the degree $k\leq 2n-1$ cohomology is independent of $n$. ∎
Decomposing $\wedge^{\bullet}(V_{g}\otimes\widetilde{W}_{n})$ into irreducible representations and applying Looijenga’s result becomes rather complicated, but in the smallest nontrivial case $n=2$ we have the following:
Example 4.10.
In degree $k\leq(2g-2)/3$, $H^{k}(\mathcal{M}^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},2,1);\mathbb{Q})$ is isomorphic to
$\bigoplus_{\begin{subarray}{c}3r+m\leq k\\ r\text{ even}\end{subarray}}\bigoplus_{i=0}^{\frac{r}{2}}\Big(\bigoplus_{\lambda|2i}E_{\lambda}\otimes H^{*}(\mathcal{M}_{g};\mathbb{Q})\Big)_{(k-3r-m+2i)}\otimes(\mathbb{Q}[\alpha_{2},\beta_{4}])_{(m)}$
$\oplus$ $\bigoplus_{\begin{subarray}{c}3r+m\leq k\\ r\text{ odd}\end{subarray}}\bigoplus_{i=1}^{\frac{r+1}{2}}\Big(\bigoplus_{\lambda|2i-1}E_{\lambda}\otimes H^{*}(\mathcal{M}_{g};\mathbb{Q})\Big)_{(k-3r-m+2i-1)}\otimes(\mathbb{Q}[\alpha_{2},\beta_{4}])_{(m)},$
where $\lambda=1^{\ell_{1}}2^{\ell_{2}}3^{\ell_{3}}\cdots$ is a numerical partition of $2i$ (resp. 2i1)2i-1) and
$E_{\lambda}=\bigotimes_{t\geq 1}u_{2}^{\ell_{t}\cdot\max\{2,t-1\}}\mathbb{Q}[c_{2},c_{4},\cdots,c_{2\ell_{t}}].$
Here $(m)$ denotes the degree-$m$ component, and the subscript of each variable is its degree. Note that in this range $H^{*}(\mathcal{M}_{g};\mathbb{Q})$ is freely generated by the classes $\kappa_{i}\in H^{2i}(\mathcal{M}_{g};\mathbb{Q})$.
4.3. Stable rational cohomology of $M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d)$ and $\mathcal{M}^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d)$
Let $\pi_{G}$ denote the projection map $M_{B}^{G}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}$ for $G=\operatorname{GL},\operatorname{SL},\operatorname{PGL}$.
Proof of Theorem 1.4.
We start with $\operatorname{SL}$. Recall from the proof of Corollary 3.5 that $R^{i}\pi_{\operatorname{PGL}*}\mathbb{Q}=R^{i}\pi_{\operatorname{SL}*}\mathbb{Q}$ up to degree
$2n^{2}(1-p_{n}^{-1})(g-1)\geq n^{2}(g-1)\geq 2(g-1)(n-1)+2$
(since the statement of the theorem is only meaningful when $g\geq 2$ and $n\geq 2$). The local systems $R^{i}\pi_{\operatorname{PGL}*}\mathbb{Q}$ are pulled back from those described in Theorem 1.1, so the description of the cohomology in degrees $\leq 2(g-1)(n-1)+2$ is exactly as before, except now all local systems are pulled back to $\mathcal{M}_{g,1}.$
Now we consider $\operatorname{GL}$. Combining equations (13) and (17) we have $R\pi_{\operatorname{GL}*}\mathbb{Q}=R\pi_{\operatorname{PGL}*}\mathbb{Q}\otimes R\pi_{\operatorname{GL}_{1}*}\mathbb{Q}$, where $R^{i}\pi_{\operatorname{PGL}*}\mathbb{Q}$ is as described before and $R^{i}\pi_{\operatorname{GL}_{1}*}\mathbb{Q}=\wedge^{i}V_{g}$ by Proposition 4.9. We conclude that
$H^{*}(M_{B}^{\operatorname{GL}}(\mathcal{C}_{g,1},n,d);\mathbb{Q})\cong H^{*}(\mathcal{M}_{g,1},\textstyle{\bigwedge^{\bullet}}(V_{g}\otimes\widetilde{W}_{n})\otimes\operatorname{Sym}^{\bullet}(V_{\circ}\otimes\widetilde{W}_{n})\otimes\textstyle{\bigwedge^{\bullet}}V_{g}),$
matching the expression shown once we identify $(V_{g}\otimes\widetilde{W}_{n})\oplus V_{g}=V_{g}\otimes W_{n}$, where $W_{n}$ is defined in (18). Finally, the stabilization argument proceeds exactly as before, except that the needed input is now $H^{*}(\mathcal{M}_{g,1},S_{\langle\lambda\rangle}(V_{g}))$, which is deduced from [26] and [44, Appendix B]. ∎
As before, we present example computations in the case $n=2$.
Example 4.11.
In degree $k\leq(2g-2)/3$, $H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},2,1);\mathbb{Q})$ is isomorphic to
$\bigoplus_{\begin{subarray}{c}3r+m\leq k\\ r\text{ even}\end{subarray}}\bigoplus_{i=0}^{\frac{r}{2}}\big(\bigoplus_{\lambda|2i}\widetilde{E}_{\lambda}\otimes H^{*}(\mathcal{M}_{g,1};\mathbb{Q})\big)_{(k-3r-m+2i)}\otimes(\mathbb{Q}[\alpha_{2},\beta_{4}])_{(m)}$
$\oplus$ $\bigoplus_{\begin{subarray}{c}3r+m\leq k\\ r\text{ odd}\end{subarray}}\bigoplus_{i=1}^{\frac{r+1}{2}}\big(\bigoplus_{\lambda|2i-1}\widetilde{E}_{\lambda}\otimes H^{*}(\mathcal{M}_{g,1};\mathbb{Q})\big)_{(k-3r-m+2i-1)}\otimes(\mathbb{Q}[\alpha_{2},\beta_{4}])_{(m)},$
with notation as in Example 4.10, except that now for a numerical partition $\lambda=1^{\ell_{1}}2^{\ell_{2}}3^{\ell_{3}}\cdots$ we have
$\widetilde{E}_{\lambda}=\bigotimes_{t\geq 1}u_{2}^{\ell_{t}\cdot(t-1+\sigma(t))}\mathbb{Q}[c_{2},c_{4},\cdots,c_{2\ell_{t}}],$
with $\sigma(t)=1$ if $t=1$ or $2$, and $\sigma(t)=0$ otherwise. In this range $H^{*}(\mathcal{M}_{g,1};\mathbb{Q})\cong H^{*}(\mathcal{M}_{g};\mathbb{Q})[e_{2}]$.
In the cases of $\operatorname{PGL}_{2}$ and $\operatorname{SL}_{2}$, we only needed to decompose representations of the form $\wedge^{r}V_{g}$, which can be done as described in (20). However, in the case of $\operatorname{GL}_{2}$, there are also representations of the form $S_{\langle 1^{i}\rangle}(V_{g})\otimes S_{\langle 1^{j}\rangle}(V_{g})$; the relevant part of the Littlewood-Richardson rule is [28, p. 510]
$S_{\langle 1^{i}\rangle}(V_{g})\otimes S_{\langle 1^{j}\rangle}(V_{g})=\bigoplus\limits_{a=0}^{min(i,j)}\bigoplus\limits_{b=0}^{a}S_{\langle 2^{a-b}1^{i+j-2a}\rangle}(V_{g}).$
Example 4.12.
In degree $k\leq(2g-2)/3$, $H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},2,1);\mathbb{Q})$ is isomorphic to
$\bigoplus_{r_{1}+3r_{2}+m\leq k}\bigoplus_{\begin{subarray}{c}0\leq i\leq r_{1},r_{1}-i\in 2\mathbb{Z}\\ 0\leq j\leq r_{2},r_{2}-j\in 2\mathbb{Z}\end{subarray}}\bigoplus_{a=0}^{\min(i,j)}\bigoplus_{b=0}^{a}H^{k-m-r_{1}-3r_{2}}\big(\mathcal{M}_{g,1};S_{\langle 2^{a-b}1^{i+j-2a}\rangle}(V_{g})\big)\otimes(\mathbb{Q}[\alpha_{2},\beta_{2}])_{(m)},$
where $H^{*}(\mathcal{M}_{g,1};S_{\langle 2^{a-b}1^{i+j-2a}\rangle}(V_{g}))$ can be deduced from [44, Appendix B].
Appendix A Stable intersection cohomology beyond the smooth case (by Anne Larsen and Mirko Mauri)
The goal of this appendix is to extend Theorems 1.1 and 1.4 beyond the smooth case $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, to arbitrary degrees $d$. More precisely, we prove that for any $d\in\mathbb{Z}/n\mathbb{Z}$, the intersection cohomology of the universal character variety stabilizes to a value independent of $d$. This is a relative version of the degree-independence of the stable intersection cohomology of Dolbeault moduli spaces on a single curve [38, Proposition 1.9].
Theorem A.1.
For any $d\in\mathbb{Z}/n\mathbb{Z}$ and $g\geq 2$, the intersection cohomology groups
$I\!H^{k}(M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d),\mathbb{Q}),I\!H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q}),I\!H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})$
are independent of the degree $d$ for k<4(g1)(n1)2k<4(g-1)(n-1)-2. In particular, they are isomorphic to the groups $H^{k}(M^{G}_{B}(\mathcal{C}_{g,*},n,d),\mathbb{Q})$ for $d\in(\mathbb{Z}/n\mathbb{Z})^{*}$, computed in Theorems 1.1 and 1.4.
The bound 4(g1)(n1)24(g-1)(n-1)-2 is optimal, except for $n=2$; see §A.2.3.
Notation. Given $Z\subset X$ a closed and irreducible subvariety of a complex algebraic variety $X$, or more generally a closed embedding of integral Deligne–Mumford stacks, and $\mathcal{L}$ a local system on an open subset $U\subset Z$, we write $IC_{Z}(\mathcal{L})$ for the intermediate extension of $\mathcal{L}$. We adopt the topologist’s convention: $IC_{Z}(\mathcal{L})$ is a bounded complex with trivial cohomology sheaves in negative degrees, and $IC_{Z}(\mathcal{L})[\dim Z]$ is a perverse sheaf. The intersection complex $IC_{X}\coloneqq IC_{X}(\mathbb{Q}_{X})$ is the intermediate extension of the constant sheaf, and its cohomology is the intersection cohomology with middle perversity and rational coefficients, denoted $I\!H^{*}(X)\coloneqq H^{*}(X,IC_{X})$. The derived category of constructible complexes of sheaves on $X$ with rational coefficients is denoted $D^{b}_{c}(X,\mathbb{Q})$. (Definitions can be found in [7, §2].)
A.1. The direct image of the intersection complex along the Hitchin fibration
A.1.1. Spectral curves
Fix $g\geq 2$. For any curve $C]\in\mathcal{M}_{g}$, consider the affine spaces
$\mathbb{A}_{0}(C,n)\coloneqq\bigoplus_{i=2}^{n}H^{0}(C,\omega_{C}^{\otimes i})\subset\bigoplus_{i=1}^{n}H^{0}(C,\omega_{C}^{\otimes i})\eqqcolon\mathbb{A}(C,n).$
Each point $a=(a_{i})_{i}\in\mathbb{A}(C,n)$ defines the spectral curve
$C^{\operatorname{spec}}_{a}\coloneqq V(t^{n}+(\pi^{*}a_{1})t^{n-1}+(\pi^{*}a_{2})t^{n-2}+\ldots+\pi^{*}a_{n})\subset T^{*}C$
of degree $n$ over $C$, where $t\in H^{0}(T^{*}C,\pi^{*}\omega_{C})$ is the tautological section of the line bundle $\pi^{*}\omega_{C}$, i.e., the pullback of the canonical line bundle of $C$ along the projection $\pi\colon T^{*}C\to C$. This description allows us to define various loci in $\mathbb{A}_{0}(C,n)$ and $\mathbb{A}(C,n)$, in particular:
Definition A.2.
The smooth (resp. elliptic) locus of $\mathbb{A}_{0}(C,n)$ or $\mathbb{A}(C,n)$ is the open subset over which the family of spectral curves is smooth (resp. integral).
Recall the (proper) Hitchin morphisms
$\chi^{\operatorname{GL}}\colon M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)\to\mathbb{A}(C,n),\,\,\chi^{\operatorname{SL}}\colon M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\to\mathbb{A}_{0}(C,n),$
$\chi^{\operatorname{PGL}}\colon M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)\to\mathbb{A}_{0}(C,n),$
which assign to each Higgs bundle $(E,\theta)$ the characteristic polynomial of its Higgs field $\theta$. The fibers of these morphisms are related to the spectral curves as follows:
  • $(\chi^{\operatorname{GL}})^{-1}(a)$ is isomorphic to the compactified Jacobian $\overline{\mathrm{Jac}}^{d}(C^{\operatorname{spec}}_{a})$ parametrizing rank 1 torsion free sheaves on $C^{\operatorname{spec}}_{a}$ of degree $d+(n^{2}-n)(g-1)$, semistable with respect to the canonical polarization;
  • $(\chi^{\operatorname{SL}})^{-1}(a)$ is isomorphic to $\overline{\operatorname{Prym}}^{d}(C^{\operatorname{spec}}_{a}/C)$, i.e., the fiber over $L$ of the norm map $\overline{\mathrm{Jac}}^{d}(C^{\operatorname{spec}}_{a})\to\mathrm{Jac}^{d}(C)$ given by $\mathcal{I}\mapsto\det(\pi_{*}\mathcal{I})$;
  • $(\chi^{\operatorname{PGL}})^{-1}(a)$ is isomorphic to the quotient $\overline{\operatorname{Prym}}^{d}(C^{\operatorname{spec}}_{a}/C)/\mathrm{Jac}^{0}(C)[n]$, where the action of $\mathrm{Jac}^{0}(C)[n]$ is by tensor product with the pullback bundle on $C_{a}^{\operatorname{spec}}$.
See [24, §5.1] (over the smooth locus), [1, Proposition 3.4] (over the elliptic locus), and [46] and [17, Proposition 6.1] (in general).
The symmetries of the moduli spaces force tight relations between the direct images of the intersection complexes of $M^{G}_{\mathrm{Dol}}(C,n)$ along the maps $\chi^{G}$ for $G=\operatorname{GL},\operatorname{SL},\operatorname{PGL}$, which we explore in the following. We first compare $\chi^{\operatorname{SL}}$ and $\chi^{\operatorname{PGL}}$.
A.1.2. $\operatorname{SL}$ vs $\operatorname{PGL}$
The group $\Gamma\coloneqq\mathrm{Jac}^{0}(C)[n]$ acts on $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ by tensor product, as follows: $L_{\gamma}\cdot(E,\theta)=(E\otimes L_{\gamma},\theta)$ for any $L_{\gamma}\in\Gamma$. The Hitchin morphism $\chi^{\operatorname{SL}}$ is $\Gamma$-invariant. The direct image $R\chi_{*}^{\operatorname{SL}}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$ inherits a $\Gamma$-action and decomposes in the sum of the $\Gamma$-invariant and the $\Gamma$-variant (i.e., non-invariant) parts:
$R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}=(R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}\oplus(R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\mathrm{var}}.$ (22)
Since the $\Gamma$-quotient of $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is $M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)$, the well-known Lemma A.4 gives the following characterization of the invariant part.
Proposition A.3.
There is an isomorphism in $D^{b}_{c}(\mathbb{A}_{0}(C,n),\mathbb{Q})$
$(R\chi_{*}^{\operatorname{SL}}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}\cong R\chi_{*}^{\operatorname{PGL}}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)}.$
Lemma A.4.
Let $X$ be an irreducible complex variety acted on by a finite group $\Gamma$, and let $f\colon X\to X/\Gamma\eqqcolon Y$ be the geometric quotient. Then
$(f_{*}IC_{X})^{\Gamma}\cong IC_{Y}.$
Proof.
Let $U\subset Y$ be a dense smooth open subset over which $f$ is étale. Since $f$ is finite, $f_{*}IC_{X}$ is an intersection complex, as it satisfies the support and cosupport conditions for intersection complexes [7, (12), (13)]. This means that
$f_{*}IC_{X}\cong IC_{Y}(f_{*}\mathbb{Q}_{f^{-1}(U)}).$ (23)
The quotient endows these complexes with a natural $\Gamma$-action, intertwined by the isomorphism (23). We conclude that $(f_{*}IC_{X})^{\Gamma}=IC_{Y}((f_{*}\mathbb{Q}_{f^{-1}(U)})^{\Gamma})=IC_{Y}(\mathbb{Q}_{U})$. ∎
The invariant part has no proper direct summand on the elliptic locus.
Proposition A.5.
The complex $R\chi_{*}^{\operatorname{PGL}}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)}$ has full support on the elliptic locus. In particular, the codimension of each proper support is $\geq 2(g-1)(n-1)-1$.
Proof.
By [40, §7], a Ngô fibration with integral fibers (in particular, the Hitchin fibration over the elliptic locus) has full support. Alternatively, see [38, Theorem 3.13] together with Proposition A.10, or [6] together with Proposition A.10 and Proposition A.6. ∎
Proposition A.6.
Let $e\in(\mathbb{Z}/n\mathbb{Z})^{*}$ and $L^{\prime}\in\mathrm{Pic}^{e}(C)$. Then
$R\chi_{*}^{\operatorname{GL}}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)},\quad R\chi_{*}^{\operatorname{PGL}}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)},\quad R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$
are respectively direct summands of
$R\chi_{*}^{\operatorname{GL}}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,e)},\quad R\chi_{*}^{\operatorname{PGL}}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,e)},\quad R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L^{\prime})}.$
Proof.
This follows from the degree-independence of BPS sheaves [27] in the case of $G=\operatorname{GL}$ by [38, Proposition 3.9], thus also for $G=\operatorname{PGL}$ by Proposition A.10. The same argument works for $G=\operatorname{SL}$ mutatis mutandis; cf. [42, Lemma 8.4]. Note that $R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$ is independent of $L\in\mathrm{Pic}^{d}(C)$. ∎
The variant part instead is supported on the endoscopic locus, which we now define.
Definition A.7.
The endoscopic locus $\mathbb{A}_{0}(C,n)^{\operatorname{end}}$ is the locus of points $a\in\mathbb{A}_{0}(C,n)$ such that the group scheme $\operatorname{Prym}^{0}(C^{\operatorname{spec}}_{a}/C)$ is disconnected, where $\operatorname{Prym}^{0}(C^{\operatorname{spec}}_{a}/C)$ is the kernel of the norm map $\mathrm{Jac}^{0}(C^{\operatorname{spec}}_{a})\to\mathrm{Jac}^{0}(C)$ given by $\mathcal{L}\mapsto\det(\pi_{*}\mathcal{L})\otimes\det(\pi_{*}\mathcal{O}_{C^{\operatorname{spec}}_{a}})^{\vee}$.
Proposition A.8.
The complex $(R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\mathrm{var}}$ is supported on the endoscopic locus. In particular, its support has codimension $\geq(1-\frac{1}{p})(g-1)n^{2}$ where $p$ is the smallest prime dividing $n$.
Proof.
By Proposition A.6, it suffices to consider the case $\deg(L)\in(\mathbb{Z}/n\mathbb{Z})^{*}$, so that $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is smooth, and $IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}=\mathbb{Q}_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$. Now, let $IC_{Z}(\mathcal{L})[-k]$ be a simple proper direct summand of $(R\chi^{\operatorname{SL}}_{*}\mathbb{Q}_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\mathrm{var}}$. Suppose by contradiction that $Z$ does not lie in the endoscopic locus, and choose a general point $a\in Z$. Then $\mathcal{L}_{a}$ is a $\Gamma$-equivariant direct summand of $(R^{k}\chi^{\operatorname{SL}}_{*}\mathbb{Q}_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})_{a}=H^{k}((\chi^{\operatorname{SL}})^{-1}(a),\mathbb{Q})$, endowed with a non-trivial $\Gamma$-action if $\mathcal{L}_{a}\neq 0$. Since $a$ does not lie in the endoscopic locus, the action of $\Gamma$ on the fiber $(\chi^{\operatorname{SL}})^{-1}(a)$ extends to the action of the connected group $\operatorname{Prym}^{0}(C_{a}^{spec})$, so $\Gamma$ acts trivially on the fiberwise cohomology groups $H^{*}((\chi^{\operatorname{SL}})^{-1}(a),\mathbb{Q})$. We conclude that $\mathcal{L}=0$, i.e., the variant part of $R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$ is supported on the endoscopic locus. Finally, the codimension of the endoscopic locus is $(1-\frac{1}{p})(g-1)n^{2}$ by [17, Corollary 1.3, Lemma 7.1]. ∎
Remark A.9.
The reduction to the smooth case in the proof of Proposition A.8 requires the degree-independence of BPS sheaves (Proposition A.6). Alternatively, one may observe that the integer $k$ in the previous proof can be taken to be $2c$, where $c$ is the relative dimension of $\chi^{\operatorname{SL}}$, by Ngô freeness and $\delta$-regularity; see [35, Proposition 1.5] and [4, Theorem A]. In top degree, the direct images of the intersection complex and constant sheaf coincide by [38, Proposition 3.3], i.e., $R^{2c}\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}\cong R^{2c}\chi^{\operatorname{SL}}_{*}\mathbb{Q}_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}$, which makes the proof work in any degree. For our purposes, deep arguments using the degree-independence or Ngô freeness can be avoided: it suffices to prove the proposition over the elliptic locus, over which $M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)$ is smooth in any degree. The codimension estimate of the complement of the nonendoscopic elliptic locus is worse, namely $\geq 2(g-1)(n-1)-1$, but sufficient for Theorem A.1.
A.1.3. $\operatorname{GL}$ vs $\operatorname{PGL}$
Recall that $H^{0}(C,\omega_{C})$ acts on $\mathbb{A}(C,n)$ via the translation action of $H^{0}(C,\omega_{C})\xrightarrow{\pi^{*}}H^{0}(T^{*}C,\pi^{*}\omega_{C})$ on the tautological section $t\in H^{0}(T^{*}C,\pi^{*}\omega_{C})$. Consider the quotient map $q\colon\mathbb{A}(C,n)\cong H^{0}(C,\omega_{C})\oplus\mathbb{A}_{0}(C,n)\to\mathbb{A}_{0}(C,n)$ given explicitly by
$p(t)=t^{n}+\sum^{n}_{i=1}\pi^{*}a_{i}t^{n-1}\mapsto p(t-\pi^{*}a_{1}/n)\eqqcolon t^{n}+\sum^{n}_{i=2}\pi^{*}b_{i}t^{n-1}.$
Proposition A.10.
There is an isomorphism in $D^{b}_{c}(\mathbb{A}(C,n),\mathbb{Q})$
$R\chi_{*}^{\operatorname{GL}}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)}\cong H^{*}(\mathrm{Pic}^{0}(C),\mathbb{Q})\otimes q^{*}R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)}.$
Proof.
Recall that there is a commutative diagram
${M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)\times M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{/\,\Gamma}$$\scriptstyle{(\chi^{\mathbb{G}_{m}},\chi^{\operatorname{SL}})}$${M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi^{\operatorname{GL}}}$${H^{0}(C,\omega_{C})\times\mathbb{A}_{0}(C,n)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cong}$${\mathbb{A}(C,n)}$
where the top horizontal arrow is given by $((E,\theta),(F,\psi))\mapsto(E\otimes F,\theta+\psi/n\otimes\mathrm{id}_{F})$, and the bottom horizontal arrow by $(a_{1},p(t))\mapsto p(t+\pi^{*}a_{1}/n)$. In fact, the top horizontal arrow is the quotient by the free $\Gamma$-action given by
$L_{\gamma}\cdot((E,\theta),(F,\psi))=((E\otimes L_{\gamma},\theta\otimes\mathrm{id}),(F\otimes L_{\gamma}^{-1},\psi\otimes\mathrm{id}))$
for any $L_{\gamma}\in\Gamma$. By Lemma A.4 and the Künneth formula in intersection cohomology [3, Proposition 2], we obtain
$R\chi^{\operatorname{GL}}_{*}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,n,d)}$ $\cong(R(\chi^{\mathbb{G}_{m}},\chi^{\operatorname{SL}})_{*}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)\times M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}$
$\cong(R\chi^{\mathbb{G}_{m}}_{*}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)}\boxtimes R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}.$
Since the Hitchin map $\chi^{\mathbb{G}_{m}}\colon M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)\to H^{0}(C,\omega_{C})$ is isomorphic to the projection $\mathrm{Pic}^{0}(C)\times H^{0}(C,\omega_{C})\to H^{0}(C,\omega_{C})$, we have
$R\chi^{\mathbb{G}_{m}}_{*}IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,0)}\cong\mathbb{Q}_{H^{0}(C,\omega_{C})}\otimes H^{*}(\mathrm{Pic}^{0}(C),\mathbb{Q}).$
Further, since the translation action of $\Gamma$ on $\mathrm{Pic}^{0}(C)$ induces the trivial action in cohomology, we obtain
$(R\chi^{\mathbb{G}_{m}}_{*}$ $IC_{M^{\operatorname{GL}}_{\mathrm{Dol}}(C,1,d)}\boxtimes R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}$
$\cong\big((\mathbb{Q}_{H^{0}(C,\omega_{C})}\otimes H^{*}(\mathrm{Pic}^{0}(C),\mathbb{Q})\boxtimes R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)}\big)^{\Gamma}$
$\cong H^{*}(\mathrm{Pic}^{0}(C),\mathbb{Q})\otimes q^{*}(R\chi^{\operatorname{SL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)})^{\Gamma}$
$\cong H^{*}(\mathrm{Pic}^{0}(C),\mathbb{Q})\otimes q^{*}R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L)},$
where the last isomorphism follows from Proposition A.3. ∎
A.1.4. Shifts of summands and codimension of supports
Lemma A.11.
Let $f\colon X\to Y$ be a proper morphism of algebraic varieties satisfying $R^{>2c}f_{*}IC_{X}=0$ for $c:=\dim X-\dim Y$. For all direct summands $IC_{Z}(\mathcal{L})[-k]$ of $Rf_{*}IC_{X}$ with $Z\subsetneq Y$, we have
$k\geq 2\operatorname{codim}Z.$
Proof.
If $IC_{Z}(\mathcal{L})[-k]$ is a direct summand of $Rf_{*}IC_{X}$, then $IC_{Z}(\mathcal{L})[\dim Z][c-k+\operatorname{codim}Z]$ is a direct summand of $Rf_{*}IC_{X}[\dim X]$. By the properness of $f$ and since $IC_{X}[\dim X]$ is self-dual, Poincaré-Verdier duality gives
$^{\mathfrak{p}}\mathcal{H}^{c-k+\operatorname{codim}Z}(Rf_{*}IC_{X}[\dim X])\cong{}^{\mathfrak{p}}\mathcal{H}^{-(c-k+\operatorname{codim}Z)}(Rf_{*}IC_{X}[\dim X])^{\vee},$
which implies that $IC_{Z}(\mathcal{L}^{\vee})[-k-2(c-k+\operatorname{codim}Z)]$ is also a direct summand of $Rf_{*}IC_{X}$. Since $R^{>2c}f_{*}IC_{X}=0$, we must have $k+2(c-k+\operatorname{codim}Z)\leq 2c$, i.e., $k\geq 2\operatorname{codim}Z$. ∎
Remark A.12.
The cohomological bound $R^{>2c}f_{*}IC_{X}=0$ holds if $X$ admits a Whitney stratification $X=\bigsqcup X_{i}$ such that for any $y\in Y$
$\operatorname{codim}_{X}(X_{i})=2\operatorname{codim}_{f^{-1}(y)}(X_{i}\cap f^{-1}(y));$
see [38, Proposition 3.3]. This is the case if $f$ is a Lagrangian fibration from a possibly singular symplectic variety [38, Proposition 3.6]. In particular, it holds for the Hitchin fibration $\chi^{G}(C)\colon M^{G}_{\mathrm{Dol}}(C)\to\mathbb{A}(C,G)$ by [4, Proposition 1.5].
A.2. Degree-independence of the stable intersection cohomology of the universal character variety
A.2.1. The case of $\operatorname{PGL}$
Recall the relative Hitchin fibration
$\chi^{\operatorname{PGL}}\colon M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathbb{A}_{0}(\mathcal{C}_{g},n),$
where $\mathbb{A}_{0}(\mathcal{C}_{g},n)\to\mathcal{M}_{g}$ is the affine bundle with fiber $\mathbb{A}_{0}(C,n)$ over $[C]$. Consider the universal spectral curve $\mathscr{C}^{\operatorname{spec}}\to\mathbb{A}_{0}(\mathcal{C}_{g},n)$ over $\mathcal{C}_{g}\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n)$. Let $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$ be the locus over which $\mathscr{C}^{\operatorname{spec}}$ is smooth.
Proposition A.13.
The complex $R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}|_{\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}}$ has full support and is independent of $d$.
Proof.
Let $\mathrm{Pic}^{0}(\mathcal{C}_{g}\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n)/\mathbb{A}_{0}(\mathcal{C}_{g},n))$ (resp. $\mathrm{Pic}^{0}(\mathscr{C}^{\operatorname{spec}}/\mathbb{A}_{0}(\mathcal{C}_{g},n))$) be the Picard functor of the pullback of the universal curve $\mathcal{C}_{g}\to\mathcal{M}_{g}$ (resp. of the universal spectral curve) over $\mathbb{A}_{0}(\mathcal{C}_{g},n)$. The kernel of the norm map
$\mathrm{Pic}^{0}(\mathscr{C}^{\operatorname{spec}}/\mathbb{A}_{0}(\mathcal{C}_{g},n))$ $\to\mathrm{Pic}^{0}(\mathcal{C}_{g}\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n)/\mathbb{A}_{0}(\mathcal{C}_{g},n))$
$\mathcal{L}$ $\mapsto\det(\pi_{*}\mathcal{L})\otimes\det(\pi_{*}\mathcal{O}_{\mathscr{C}^{\operatorname{spec}}})^{\vee}$
is the group stack denoted $\operatorname{Prym}^{0}(\mathscr{C}^{\operatorname{spec}}/\mathcal{C}_{g}\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n))$. Define the group stack
$\mathscr{P}\coloneqq\operatorname{Prym}^{0}(\mathscr{C}^{\operatorname{spec}}/\mathcal{C}_{g}\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n))/\Gamma(\mathcal{C}_{g}),$ (24)
where $\Gamma(\mathcal{C}_{g})$ is the $n$-th torsion of the universal Picard functor over $\mathcal{M}_{g}$, pulled back over $\mathbb{A}_{0}(\mathcal{C}_{g},n)$, i.e., $\mathrm{Pic}^{0}(\mathcal{C}_{g}/\mathcal{M}_{g})[n]\times_{\mathcal{M}_{g}}\mathbb{A}_{0}(\mathcal{C}_{g},n)$. Note that $M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)$ is a (smooth) $\mathscr{P}$-torsor over $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$. Then over $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$ the intersection complex $IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}$ is the constant sheaf $\mathbb{Q}$, and by [5, Lemma 1.3.5] we see that the restrictions of $R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}$ and $R\chi_{*}\mathbb{Q}_{\mathscr{P}}$ to $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{0}$ are isomorphic. The latter is independent of $d$. ∎
Therefore, the only part of $R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}$ that depends on $d$ is supported on the complement of $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$.
Proposition A.14.
If $IC_{Z}(\mathcal{L})[-k]$ is a direct summand of $R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}$ for some $Z\subsetneq\mathbb{A}_{0}(\mathcal{C}_{g},n)$, then we have $k\geq 4(g-1)(n-1)-2$.
Proof.
Choose $C]\in\mathcal{M}_{g}$ such that $Z\cap\mathbb{A}_{0}(C,n)\neq\emptyset$. By construction, the relative Betti moduli space $M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ is a locally trivial family, and so the relative Dolbeault moduli space $M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)\to\mathcal{M}_{g}$ is too by Simpson’s relative nonabelian Hodge correspondence [49, Theorem 9.11, Lemma 9.14]. The local triviality implies that
$IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d)}|_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)}\cong IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d)}.$
Therefore, $IC_{Z}(\mathcal{L})[-k]|_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C,n,d))}$ is a direct summand of
$R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d))}|_{\mathbb{A}_{0}(C,n)}\cong R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(C_{g},n,d))}.$
Note that $Z\cap\mathbb{A}_{0}(C,n)\subsetneq\mathbb{A}_{0}(C,n)$, since $Z$ avoids $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$ by Proposition A.13, and $\mathbb{A}_{0}(\mathcal{C}_{g},n)^{\circ}$ is dense in each $\mathbb{A}_{0}(C,n)$. By Proposition A.5,
$\operatorname{codim}_{\mathbb{A}_{0}(C,n)}(Z\cap\mathbb{A}_{0}(C,n))\geq 2(g-1)(n-1)-1$
for all $C]\in\mathcal{M}_{g}$ such that $Z\cap\mathbb{A}_{0}(C,n)\neq\emptyset$. In particular, we have
$\operatorname{codim}_{\mathbb{A}_{0}(\mathcal{C}_{g},n)}(Z)\geq 2(g-1)(n-1)-1$
and so the result follows from Lemma A.11 and Remark A.12.
Proposition A.15.
The intersection cohomology
$I\!H^{k}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d))\cong I\!H^{k}(M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g},n,d))$
is independent of $d$ for k<4(g1)(n1)2k<4(g-1)(n-1)-2.
Proof.
The direct image $R\chi_{*}^{\operatorname{PGL}}IC_{M_{\mathrm{Dol}}^{\operatorname{PGL}}(\mathcal{C}_{g},n,d)}$ decomposes into a sum of simple shifted perverse sheaves fully supported on $\mathbb{A}_{0}(\mathcal{C}_{g},n)$ and a sum of those supported on proper subvarieties of $\mathbb{A}_{0}(\mathcal{C}_{g},n)$. By Proposition A.13 the first term is independent of $d$. By Proposition A.14 the cohomology of the second term vanishes in degree <4(g1)(n1)2<4(g-1)(n-1)-2. So $I\!H^{k}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,d),\mathbb{Q})$ is independent of $d$ for k<4(g1)(n1)2k<4(g-1)(n-1)-2. ∎
A.2.2. From $\operatorname{PGL}$ to $\operatorname{GL}$ and $\operatorname{SL}$
Lacking a relative nonabelian Hodge correspondence for degree $d\neq 0$ and $G=\operatorname{GL}$ or $\operatorname{SL}$, we reduce Theorem A.1 for these groups to the corresponding statement for $\operatorname{PGL}$ via the fiberwise nonabelian Hodge correspondence. Let $\pi_{G}\colon M^{G}(\mathcal{C}_{g,1},n,d)\to\mathcal{M}_{g,1}$ be the projection map, and set $\mathcal{K}^{G}_{B}\coloneqq R\pi_{G,*}IC_{M^{G}_{\mathrm{Dol}}(\mathcal{C}_{g,1},n,d)}$.
Proposition A.16.
Set $M\coloneqq(1-\frac{1}{p})(2g-2)n^{2}$, where $p$ is the smallest prime dividing $n$, and $N\coloneqq 4(g-1)(n-1)-2$. There is an isomorphism
$I\!H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})\cong I\!H^{k}(M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})$
for $k<M$. Furthermore, these groups are independent of $d$ for $k<N$.
Proof.
Since the morphism $f\colon M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d)\to M_{B}^{\operatorname{PGL}}(\mathcal{C}_{g,1},n,d)$ is a quotient by the relative group object $\mathcal{T}_{g,1}\times_{\mathcal{M}_{g,1}}\mathrm{Hom}(\pi_{1}(S_{g}),\mu_{n})$, we write, as in (22) and Lemma A.3, that
$\mathcal{K}^{\operatorname{SL}}_{B}\cong(\mathcal{K}^{\operatorname{SL}}_{B})^{\Gamma}\oplus(\mathcal{K}^{\operatorname{SL}}_{B})^{\mathrm{var}}\cong\mathcal{K}^{\operatorname{PGL}}_{B}\oplus(\mathcal{K}^{\operatorname{SL}}_{B})^{\mathrm{var}}.$ (25)
The local triviality of $\pi_{G}$ implies that
$(\mathcal{K}^{\operatorname{SL}}_{B})^{\mathrm{var}}|_{[C]}\cong I\!H^{*}(M^{\operatorname{SL}}_{B}(S_{g,1},n,d))^{\mathrm{var}}.$
By the fiberwise ($\Gamma$-equivariant) nonabelian Hodge correspondence for $\operatorname{SL}$ [22, Theorem 4.3] (cf. [23, Remark 5 after Theorem 1]), there is an isomorphism
$I\!H^{k}(M^{\operatorname{SL}}_{B}(S_{g,1},n,d))^{\mathrm{var}}\cong I\!H^{k}(M^{\operatorname{SL}}_{\mathrm{Dol}}(C,n,L))^{\mathrm{var}}.$
This group vanishes for $k<M$ by Proposition A.8 and Lemma A.11. Hence, taking global cohomology of (25), we conclude that $I\!H^{k}(M^{\operatorname{SL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})\cong I\!H^{k}(M^{\operatorname{PGL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})$ for $k<M$. As in Proposition A.15, these intersection cohomology groups are independent of $d$ for $k<N$. ∎
Proposition A.17.
The intersection cohomology $I\!H^{k}(M^{\operatorname{GL}}_{B}(\mathcal{C}_{g,1},n,d),\mathbb{Q})$ is independent of $d$ for k<4(g1)(n1)2k<4(g-1)(n-1)-2.
Proof.
By the argument of (17) and Lemma A.4, we have $\mathcal{K}^{\operatorname{GL}}_{B}\cong\mathcal{K}^{\operatorname{PGL}}_{B}\otimes\mathcal{K}^{\mathbb{G}_{m}}_{B}$. Let $V_{g}$ be the shifted local system on $\mathcal{M}_{g,1}$ corresponding to the standard symplectic representation of the mapping class group in cohomological degree 1 in $D^{b}_{c}(\mathcal{M}_{g,1},\mathbb{Q})$. Since $\mathcal{K}^{\mathbb{G}_{m}}_{B}\cong\bigoplus_{i}\Lambda^{i}V_{g}$ is independent of $d$, the result follows from Proposition A.15. ∎
Proof of Theorem A.1.
This is the content of Propositions A.15 (for $\operatorname{PGL}$), A.16 (for $\operatorname{SL}$), and A.17 (for $\operatorname{GL}$). ∎
A.2.3. Sharpness of the bounds
We discuss the sharpness of Propositions A.14 and A.15.
Proposition A.18.
The bound $N\coloneqq 4(g-1)(n-1)-2$ is sharp for Proposition A.14 for arbitrary values of the pair $(n,g)$.
Proof.
Consider the inclusion $i\colon Z^{\circ}\hookrightarrow\mathbb{A}_{0}(\mathcal{C}_{g},n)$ in $\mathbb{A}_{0}(\mathcal{C}_{g},n)$ of the locus of nodal spectral curves which are union of two irreducible components, one of which is isomorphic to the base curve $\mathcal{C}_{g}$, i.e., such that the two components of the spectral curve have arithmetic genus $g$ and $(n-1)^{2}(g-1)+1$ respectively, and let $Z$ be its closure. By [38, Theorem 1.1], there is a direct summand of $R\chi^{\operatorname{PGL}}_{*}\mathbb{Q}_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1)}$, which is not a summand of $R\chi^{\operatorname{PGL}}_{*}IC_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,0)}$, of the form $i_{*}\mathcal{L}[-N]$ where $\mathcal{L}$ is a rank one local system on $Z^{\circ}$. ∎
Proposition A.19.
The bound $N\coloneqq 4(g-1)(n-1)-2$ in Proposition A.15 is sharp, except for $n=2$.
Proof.
Let $i_{*}\mathcal{L}[-N]$ be the direct summand of $R\chi^{\operatorname{PGL}}_{*}\mathbb{Q}_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1)}$ in the proof of Proposition A.18. By relative hard Lefschetz, there is a corresponding summand $\mathcal{L}$ of the sheaf $R^{\textrm{top}}\chi_{*}^{\operatorname{PGL}}\mathbb{Q}_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1)}|_{Z^{\circ}}$. The fiber of this sheaf at the point $a\in Z^{\circ}$ is generated by the fundamental classes of the irreducible components of $(\chi^{\operatorname{PGL}})^{-1}(a)$. The fiber $(\chi^{\operatorname{PGL}})^{-1}(a)$ is a compactified Jacobian of $C^{\operatorname{spec}}_{a}$ whose irreducible components are indexed by the degrees of the restriction of a stable line bundle on $C^{\operatorname{spec}}_{a}$ to its irreducible components; cf. [6, §4] and [37, Proposition 7.6.(ii)]. As described in the proof of Proposition A.18, the spectral curve $C^{\operatorname{spec}}_{a}$ has two irreducible components, which have (crucially) different genus when $n\neq 2$. In particular, they cannot be exchanged by the monodromy action, and thus we conclude that $R^{\mathrm{top}}\chi_{*}^{\operatorname{PGL}}\mathbb{Q}_{M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1)}|_{Z^{\circ}}$ is a trivial local system. The same must then be true of its summand $\mathcal{L}$.
Now, as $\mathcal{L}$ is of rank one, as long as $n\neq 2$, the contribution of $i_{*}\mathcal{L}[-N]=i_{*}\mathbb{Q}_{Z^{\circ}}[-N]$ to $H^{N}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1))$ is $H^{0}(Z,i_{*}\mathbb{Q}_{Z^{\circ}})=\mathbb{Q}$. We conclude that for $n\neq 2$ we have
$\dim H^{N}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,1))=\dim I\!H^{N}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},n,0))+1.$
On the other hand, if $n=2$, the monodromy of $\mathcal{L}|_{\mathbb{A}_{0}(C,n)\cap Z^{\circ}}$ is no longer trivial. The locus $Z_{C}\coloneqq Z\cap\mathbb{A}_{0}(C,2)$ is the image of the squaring map $H^{0}(C,\omega_{C})\to H^{0}(C,\omega^{\otimes 2}_{C})$, and in particular $\pi_{1}(Z^{\circ}_{C})=\mathbb{Z}/2\mathbb{Z}$. The monodromy along this loop acts on the fiber of $\mathcal{L}$ by $\mathrm{id}$; see [4, Corollary 6.20] or [37, Theorem 1.6]. Hence, if $n=2$, we have $H^{0}(Z,i_{*}\mathcal{L})=0$ and $\dim H^{N}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},2,1))=\dim I\!H^{N}(M^{\operatorname{PGL}}_{\mathrm{Dol}}(\mathcal{C}_{g},2,0)).$
Acknowledgment. Mirko Mauri was supported by Université Paris Cité and Sorbonne Université, CNRS, IMJ-PRG, F-75013 Paris, France.